### v1

# AgenticTagger: Structured Item Representation for Recommendation with LLM Agents

Abstract.
High-quality representations are a core requirement for effective recommendation. In this work, we study the problem of LLM-based descriptor generation, i.e., keyphrase-like natural language item representation generation frameworks with minimal constraints on downstream applications. We propose AgenticTagger, a framework that queries LLMs for representing items with sequences of text descriptors. However, open-ended generation provides little control over the generation space, leading to high cardinality, low-performance descriptors that renders downstream modeling challenging. To this end, AgenticTagger features two core stages: (1) a vocabulary building stage where a set of hierarchical, low-cardinality, and high-quality descriptors is identified, and (2) a vocabulary assignment stage where LLMs assign in-vocabulary descriptors to items. To effectively and efficiently ground vocabulary in the item corpus of interest, we design a multi-agent reflection mechanism where an architect LLM iteratively refines the vocabulary guided by parallelized feedback from annotator LLMs that validates the vocabulary against item data. Experiments on public and private data show AgenticTagger brings consistent improvements across diverse recommendation scenarios, including generative and term-based retrieval, ranking, and controllability-oriented, critique-based recommendation.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
Figure 1. Illustration of the proposed method. We represent items using an ordered sequence of coarse-to-fine LLM-generated natural language descriptors, directly extracting task-relevant information, such as genres for an item corpus of CDs and Vinyls.
1. Introduction
There have been numerous efforts in leveraging LLM for item representation in recommender systems. Typically, this process involves prompting LLMs for summarizations of item content (Valizadeh et al., 2025; Ren et al., 2024; Shi et al., 2025; Yada and Yamana, 2024; Wang et al., 2024c) or users’ behavioral sequence (Liu et al., 2025a, b; Jalan et al., 2024; Jiang et al., 2025; Wang et al., 2024d), which are then fed into a recommendation model as textual features. We study item-side, prompting-based discrete feature representation (i.e., natural language descriptor) generation with no constraints on the recommendation model architecture, as item representation is the foundation of effective recommendation (Rajput et al., 2023; Hou et al., 2023, 2025a, 2025b; Zhai et al., 2024; Cheng et al., 2023; Kang et al., 2021; Zhang et al., 2024b, 2025b). In particular, we seek frameworks capable of generating keyphrase-like descriptors, broadly applicable as ID or semantic features to a wide variety of downstream applications, complementary to existing efforts in building specialized recommendation frameworks exclusively for prompting LLMs for recommendations (Liu et al., 2025a, b; Shu et al., 2023; Zhao et al., 2024; Thakkar and Yadav, 2024; Wang et al., 2024b; Xia et al., 2026; Huang et al., 2025).
Figure 2. The AgenticTagger Framework. AgenticTagger exploits LLMs’ content-understanding ability to automatically construct an interpretable, hierarchical descriptor vocabulary from data. Then, relevant descriptors in the vocabulary could be assigned to corresponding items by querying LLMs in parallel over the item corpus. These features could then power various downstream applications, such as generative and term-based retrieval, ranking, and critique-based recommendation.
We argue this setting holds three advantages complementary to its alternatives. Firstly, item-side representation scales more easily for the typical case where recommender systems serve more users than their items (McAuley et al., 2015; Gao et al., 2022). Secondly, prompt-based methods align well with modern language models’ language-based interaction paradigm (Gheorghe Comanici et, 2025; OpenAI, 2024), and scales well with improvements in base LLMs without the need for fine-tuning. Thirdly, text descriptors and sparse discrete features hold irreplaceable value in recommender system, due to their flexibility in benefiting diverse recommendation architectures (Zhang et al., 2024b, 2025b; Wang et al., 2017, 2021; Kabbur et al., 2013; Cheng et al., 2016; Song et al., 2019), as well as their efficiency and scalability when deployed in keyword-based retrieval settings (Wartena et al., 2010; Google Ads Help, 2026), which complements prior studies in adjacent directions such as LLM-augmented item embeddings (Ren et al., 2024; Yada and Yamana, 2024; Valizadeh et al., 2025). Finally, as we discuss in later sections (Section 6.2), text-based descriptor representations naturally come with interpretability, and are well suited for user-facing applications.
However, despite LLMs’ strong content understanding ability, LLM generations offer no control over its choice of vocabulary, while a good set of discrete features should ideally capture meaningful frequent patterns in the item corpus (Xie et al., 2025; Rajput et al., 2023; Singh et al., 2024). In particular, as recommender system corpora generally exceed the context limit of LLMs (McAuley et al., 2015; Harper and Konstan, 2015), it is challenging to incorporate the abstract constraint of generating descriptors that meaningfully highlights its similarities and differences w.r.t. other items, as this requires a global view of the entire item corpus. For example, we find many features produced by free-form generation only occur on a single item across the corpus (Section 6.1), whereas feature-based machine learning requires meaningfully frequent patterns to converge. As we later discuss in Section 7.3), this issue leads to vocabulary explosion, where large quantities of descriptors are produced, resulting in more unique features than items, making downstream feature modeling challenging, compromising recommendation quality.
To this end, we propose AgenticTagger, a multi-LLM framework that queries LLMs for one-to-many natural language discrete features (i.e, tag-like descriptors, as shown in Figure 1) shared across items, where the descriptors come from an automatically learned low-cardinality vocabulary. Specifically, AgenticTagger features two key stages, (1) vocabulary building (Section 4.1), where AgenticTagger automatically learns a hierarchical coarse-to-fine descriptor taxonomy, drawing insights from LLM for taxonomy mining (Wan et al., 2024; Xie et al., 2025; Pham et al., 2024) and codebook learning for item representation in recommendation (Hou et al., 2023, 2025b, 2025a; Singh et al., 2024; He et al., 2025; Rajput et al., 2023); and (2) vocabulary assignment (Section 4.2), which prompts LLMs to assign relevant descriptors from the vocabulary to items, constraining their generation space. By pre-learning a set of high-quality, low-cardinality descriptors, AgenticTagger effectively constrains the LLMs’ generation space, equipping LLM with a global, hierarchical view of the item corpus, thereby improving generation quality.
Under the general AgenticTagger framework, a core design challenge is how to effectively “propagate” signals through LLM generations during the vocabulary building stage, to effectively improve the quality of the vocabulary for covering diverse items in the item corpus. To this end, following the numerous recent successes in self-improving LLM agents through verbal feedback (Shinn et al., 2023; Majumder et al., 2024; Xie et al., 2024), we adopt an LLM-based refinement loop, which refines the vocabulary by aiming for better coverage of the item corpus through LLM-generated failure reports from items that do not fit into the current vocabulary. On the other hand, unlike these prior works, AgenticTagger uniquely adopts a multi-agent architect-annotator framework, where a stronger (yet slower) architect LLM maintains the descriptor vocabulary, while numerous copies of the weaker (yet faster) annotator LLMs scan through the item corpus in parallel, making AgenticTagger easily scalable to large datasets.
We conduct experiments with AgenticTagger across public benchmarks and a private dataset, showing the broad applicability of AgenticTagger via its improvements to diverse recommendation settings including term-based retrieval, generative retrieval, and ranking with ID-like item features. We also show that these natural language descriptors natively bring interpretability and controllability to recommender systems, such as making generative recommender models compatible with critiques by controlling their decoding process via user feedback. Our contributions are as follows:
  • To the best of our knowledge, we are the first to apply agentic frameworks to feature mining to benefit recommender systems, opening up a new direction for future work.
  • We identify the core challenge in LLM-based item representation generation, namely, vocabulary mining, and propose a multi-agent framework for addressing such issue.
  • We show the value of LLM-agent-produced, ID-like natural language item representations in recommender systems, for their performance, interpretability, and flexibility.
2. Related Work
LLM-generated Features in RecSys. Prior work in introducing LLM-produced features in recommender system broadly falls into two streams: LLM-as-encoders (Zhang et al., 2024b, 2025b) and LLM-as-text-generators (Ren et al., 2024; Yada and Yamana, 2024; Valizadeh et al., 2025; Liu et al., 2024, 2025a), and this work aims at enhancing the latter. In contrast to the popular goal of generating user profiles that improve an LLM itself at recommendation task (Liu et al., 2025a, b), we consider specifically the case where the recommendation model is a general feature-aware ML model, which aligns well with most existing deployed recommendation models. Compared to the most similar works that encode LLM-produced descriptions (Ren et al., 2024; Yada and Yamana, 2024; Valizadeh et al., 2025), we are the first work to demonstrate the value of generating one-to-many discrete features, and the first work to demonstrate the value of LLM agents in feature generation for RecSys.
Discrete Representation in RecSys. Our work is also similar to a series of effort in building better discrete features for RecSys, which typically involve discretizing item representations with product quantization and its variants (Jegou et al., 2011; Ge et al., 2014; Hou et al., 2023, 2025a), RQ-VAE (van den Oord et al., 2017; Rajput et al., 2023), and k-means (Ju et al., 2025; Zhang et al., 2024c). However, these methods do not leverage the reasoning ability of LLMs, and as we will later show in the paper, discrete features produced from raw text by exploiting LLMs’ content understanding ability outperform their counterparts such as RQ-VAE-based semantic IDs.
LLM Agents for RecSys. There have been numerous recent efforts in leveraging LLM agents to improve RecSys. These frameworks generally focus on leveraging LLMs for item recommendation (Shu et al., 2023; Zhao et al., 2024; Thakkar and Yadav, 2024; Wang et al., 2024b; Xia et al., 2026; Huang et al., 2025), or enabling better interactivity by incorporating LLM and LLM agents into the recommendation serving stage (Friedman et al., 2023; He et al., 2023; Zhang et al., 2024a). However, these frameworks typically require specialized pipelines, whereas AgenticTagger focuses on feature generation, which potentially benefits diverse downstream recommendation models. Meanwhile, to the best of our knowledge, AgenticTagger is the first framework to study agentic feature generation for recommender systems.
Table 1. Summary of Notations
Notation Description
$\mathcal{U},\mathcal{I}$ Sets of users and items
$S_{u}$ Interaction history for user $u\in\mathcal{U}$
$X_{i}$ Raw attributes for item $i\in\mathcal{I}$
$\Pi,\pi$ LLMs, and a specific instance
$\mathcal{E}$ Set of natural language error feedback
AA Descriptor feature vocabulary
$a$ A specific feature descriptor ($a\in A$)
$g(\cdot)$ Transformation mapping attributes to a subset of AA
$f_{\theta}$ Recommender system parameterized by $\theta$
3. Problem Statement
Consider the standard recommendation setting with item features. Given a set of users $\mathcal{U}$ and items $\mathcal{I}$, the objective is to predict unobserved items of interest for a user $u\in\mathcal{U}$ based on their history of interactions $S_{u}\subset\mathcal{I}$. Each item $i\in\mathcal{I}$ is associated with raw textual attributes $X_{i}\in\mathcal{X}$, where $\mathcal{X}$ is the original feature space. In the context of feature generation, our goal is to find an optimal one-to-many transformation function, $g:\mathcal{X}\to\mathcal{P}(A)$, that maps these original attributes to a new discrete feature space $A$. The optimal transformation, $g^{*}$, is one that maximizes the performance of a downstream recommender system, $f$:
$g^{*}=\arg\max_{g}\mathcal{M}(f_{\theta^{*}(g)})$
where $f_{\theta^{*}(g)}$ is the recommender model trained using the generated features, and $\mathcal{M}$ is a relevant performance metric, e.g., offline metrics and key business metrics. Note that while AgenticTagger do not directly optimize for end performance due to the challenge of propagating gradients over an API-supported LLM, the best vocabulary should ideally, in our problem formulation, maximize downstream task performance. We provide an overview of the notations used in Table 1.
4. AgenticTagger
As shown in Figure 2, the overall pipeline of AgenticTagger involves two key stages: a vocabulary building stage, where we produce a descriptor vocabulary by observing typical items in the dataset, and a vocabulary assignment stage that assigns the best set of descriptors to items.
4.1. Vocabulary Building
Algorithm 1 Vocabulary Building
Input: Item corpus $\mathcal{I}$, LLM Ensemble $\Pi$
Hyperparameters: Max hierarchy depth $D_{max}$, split threshold $\tau_{split}$
Output: Hierarchical feature set $A_{all}$
$A_{all}\leftarrow\emptyset$
$p_{0}\leftarrow\text{new Feature}(\text{items}\leftarrow\mathcal{I},\text{parent}\leftarrow\text{null})$
$Q\leftarrow\{p_{0}\}$
for each $d\in\{1,\dots,D_{max}\}$ do
  if $|Q|=0$ then
   break
  end if
  $Q_{next}\leftarrow\emptyset$
  for each parent feature $p\in Q$ do
   if $|p.\text{items}|\geq\tau_{split}$ then
[…]
    $A_{all}\leftarrow A_{all}\cup A_{children}$
    $Q_{next}\leftarrow Q_{next}\cup A_{children}$
   end if
  end for
  $Q\leftarrow Q_{next}$
end for
return $A_{all}$
During the vocabulary building stage, AgenticTagger automatically learns a vocabulary from the item corpus. However, gradient-based vocabulary learning (Rajput et al., 2023; Hou et al., 2023, 2025a; Zhang et al., 2024c) is inapplicable to prompting-based LLM agents, and thus common solution is prompting an LLM in mini-batches (Wan et al., 2024; Pham et al., 2024) and refining an in-context vocabulary via LLM self-refinement (Madaan et al., 2023). However, such an iterative approach involves a non-parallelizable loop over the whole item corpus, and scales poorly to large datasets (Xie et al., 2025).
4.1.1. Hierarchical Vocabulary Construction.
In light of the above issues, in AgenticTagger, we opt to scale existing LLM-based frameworks by making the vocabulary building process parallelizable, where multiple LLMs are deployed concurrently to update a shared vocabulary that describes observed items. An overview of this process is illustrated in Algorithm 1 and Algorithm 2. Specifically, the iterative feature generation follows a hierarchical-clustering-like pipeline (Algorithm 1), where we generate an initial set of features that partitions the whole item corpus, followed by a top-down procedure to generate finer-grained features that further partition items associated with each feature into smaller subsets. For example, if the initial features includes high-level category information such as “CDs and Vinyls”, all relevant items will then form a subcluster, for generating finer-grained features such as “Rock” and “Blues”, as shown in Figure 2-(i).
4.1.2. Multi-agent Self-refinement.
At the core of this top-down generation loop is a multi-agent iterative refinement (Figure 2-(ii) and Algorithm 2) subroutine reused for each feature generation job $Q_{i}$, which aims to produce a set of features that describe common items in the corresponding subset. Our core insight is leveraging a stronger (though potentially slower) LLM, $\pi_{Architect}$, to generate and refine the vocabulary subset relevant to each $Q_{i}$, and copies of faster (and potentially weaker) LLMs, $\pi_{Annotator}$, which attempt to annotate the current item corpus using the available vocabulary and provide feedback to $\pi_{Architect}$ to improve the vocabulary. Compared to single-LLM self-refinement (Wan et al., 2024; Pham et al., 2024), the most costly $\pi_{Annotator}$ operations are parallelized, while we can still use their feedback to prompt $\pi_{Architect}$ to refine the vocabulary.
As shown in Algorithm 2, during the subroutine execution, we first prompt the $\pi_{\text{Architect}}$ LLM to generate the initial set of descriptors by observing a small subset of items $\text{Distill}(I_{sub})$ in its context, then deploy copies of (or batched queries to) the annotator LLMs $\pi_{\text{Annotator}}$ to determine if any of the currently known features apply to each item in $I_{sub}$, generating a set of error suggestions $\mathcal{E}$. The specific prompts are as shown in Section B.2. The error reports $\mathcal{E}$ then act inductive biases for prompting the vocabulary maintainer LLM $\pi_{\text{Architect}}$, to refine the current set of features. To avoid overcrowding $\pi_{\text{Architect}}$’s context, $\text{Distill}(\cdot)$ refers to a general operation that selects diverse representative samples from a larger set of item descriptions, such as K-Medoids clustering (Kaufman and Rousseeuw, 1987)’s centroids, as we use in our experiments. We repeat such an iterative refinement loop until a fixed number of optimization steps has been reached, or a desired number of items in $I_{sub}$ can be described by the generated features $A$. The union of features generated for each subset of the item corpus then forms the final feature space $A_{all}$, providing a coarse-to-fine description for the whole item corpus.
Algorithm 2 Multi-agent Self-refinement
Input: Item corpus subset $I_{sub}\subseteq\mathcal{I}$;     LLMs $\Pi=\{\pi_{\text{Architect}},\pi_{\text{Annotator}}\}$
Hyperparameters: Max refinement cycles $C_{max}$, anomaly threshold $\tau_{anom}$
Output: Descriptor vocabulary AA
$A\leftarrow\pi_{\text{Architect}}(\text{Distill}(I_{sub}))$
for $c=1$ to $C_{max}$ do
  $(I_{assigned},I_{unassigned},\mathcal{E})\leftarrow\text{ParallelAssign}(I_{sub},A,\pi_{\text{Annotator}})$
  if $|I_{unassigned}|<\tau_{anom}$ then
   break
  end if
  $A\leftarrow\pi_{\text{Architect}}(A,\text{Distill}(\mathcal{E}))$
end for
return AA
4.2. Vocabulary Assignment
Once the feature vocabulary $A_{all}$ is built, the vocabulary assignment stage for AgenticTagger is straightforward, where we simply constrain the descriptors output by the LLMs to fall into the vocabulary by prompting, i.e., $g(X_{i})=\{a\in A\mid\pi(X_{i},A)\to a\}$ for any item raw attribute $X_{i}$. For example, in AgenticTagger, for the sake of implementation simplicity, we opt to reuse the same operation from building the vocabulary, $\text{ParallelAssign}(I,A_{all},\pi_{\text{Annotator}})$, where an LLM is prompted to assign features to an item by observing all features in its context, exploiting modern LLMs’ long-context and reasoning abilities (Zhao et al., 2025). However, depending on the existing LLM serving infrastructure, this step can also be implemented as a post-processing step, where we match an LLM generation to its nearest neighbor in the entire feature vocabulary $A_{all}$, if their similarity exceeds some threshold metric (e.g., edit distance). We provide further discussion of these design choices in Section 6.1, and show the advantage of the prompting-based method, despite the longer prompts at the assignment stage.
We note that due to the prompting nature of the vocabulary assignment stage, AgenticTagger also naturally supports arbitrary natural language instructions, offering controllability from system deployers’ side. As we later showcase in our experiments (Section 5), for example, when the vocabulary itself contains hierarchical structure, one could simply instruct LLM to select one best descriptor per taxonomy level for each item, where the produced vocabulary becomes a specialized semantic ID (Rajput et al., 2023; Hou et al., 2023, 2025b, 2025a; Singh et al., 2024), where each ID is a natural language descriptor.
4.3. Complexity Analysis
4.3.1. Vocabulary Generation.
In AgenticTagger’s automated vocabulary building algorithm, the parallel calls to the annotator LLMs during the vocabulary creation phase dominate. In its most basic version, the algorithm will incur $O(C_{max}\cdot N\cdot b^{D_{max}})$ LLM calls, assuming an item will be assigned to an average of $b$ features at each depth level $d\in\{1,\dots,D_{max}\}$, which grows quickly as the depth of the feature hierarchy grows. However, the expressiveness of the feature hierarchy also grows exponentially w.r.t. the depth of the hierarchy, similar to other hierarchical discrete item representation methods (Rajput et al., 2023; Zhang et al., 2024c), and thus a few layers are often sufficient, as we later show in our experiments, 3-6 layers of features suffice good performance across public benchmarks (Table 2).
Further, since the goal of vocabulary building is solely to generate a set of feature descriptors, we can also heuristically constrain the branching factor by limiting how many features an item is assigned to at each layer, such as randomly assigning an item to one of its associated feature during the branching operation. In practice, we observe that a branching factor of 1 is also sufficient for good performance (Section 7.2), in which case the complexity w.r.t. LLM calls becomes $O(C_{max}\cdot N\cdot D_{max})$. Finally, compared to prior LLM-based document categorization methods (Wan et al., 2024; Pham et al., 2024), since AgenticTagger does not have sequential dependency between most of its dominating LLM operations (specifically, the $\pi_{\text{Annotator}}$ annotations), the overall execution time can be further reduced approximately linearly w.r.t. the number of concurrent threads in the ParallelAssign operation.
Table 2. AgenticTagger for Generative Recommendation on public benchmarks. For each metric, the best result is in bold and the second-best is underlined. Our model’s performance is reported across 3 runs. We provide additional comparison against crossing-based IDs and free-form generations in Figure 3 and Section 7.3.
Datasets Metric ID-based Feature + ID Generative
BERT4Rec SASRec FDSA S³-Rec VQ-Rec LMIndexer HSTU P5-CID TIGER AgenticTagger
Sports R@5 0.0115 0.0233 0.0182 0.0251 0.0181 0.0222 0.0258 0.0287 0.0264 0.0299 $\pm$0.0010
N@5 0.0075 0.0154 0.0122 0.0161 0.0132 0.0142 0.0165 0.0179 0.0181 0.0194 $\pm$0.0007
R@10 0.0191 0.0350 0.0288 0.0385 0.0251 0.0414 0.0426 0.0400 0.0474 $\pm$0.0002
N@10 0.0099 0.0192 0.0156 0.0204 0.0154 0.0215 0.0224 0.0225 0.0251 $\pm$0.0004
Beauty R@5 0.0203 0.0387 0.0267 0.0387 0.0434 0.0415 0.0469 0.0468 0.0454 0.0492 $\pm$0.0007
N@5 0.0124 0.0249 0.0163 0.0244 0.0311 0.0262 0.0314 0.0315 0.0321 0.0332 $\pm$0.0008
R@10 0.0347 0.0605 0.0407 0.0647 0.0741 0.0704 0.0701 0.0648 0.0728 $\pm$0.0010
N@10 0.0170 0.0318 0.0208 0.0327 0.0372 0.0389 0.0400 0.0384 0.0408 $\pm$0.0008
CDs R@5 0.0326 0.0351 0.0226 0.0213 0.0314 0.0417 0.0505 0.0492 0.0544 $\pm$0.0005
N@5 0.0201 0.0177 0.0137 0.0130 0.0209 0.0275 0.0326 0.0329 0.0358 $\pm$0.0002
R@10 0.0547 0.0619 0.0378 0.0375 0.0485 0.0638 0.0785 0.0748 0.0836 $\pm$0.0003
N@10 0.0271 0.0263 0.0186 0.0182 0.0264 0.0346 0.0416 0.0411 0.0453 $\pm$0.0002
4.3.2. Vocabulary Assignment.
Since once built, the assignment stage simply involves looping through all items of interest, the assignment loop induces $O(|\mathcal{U}|)$ LLM calls, i.e., the cost of querying $\pi_{Annotator}$ in a single-pass manner over the item corpus. To this end, once the vocabulary is built, AgenticTagger can easily scale to large item corpora with parallelized calls to LLMs. Further, since the only difference between each LLM call during the assignment stage is the item description, the inference can, in principle, be further accelerated by infrastructure optimizations such as KV-Cache (Pope et al., 2023), which we leave for future work.
4.4. Modeling AgenticTagger Features
Once the process is complete, AgenticTagger assigns a fixed maximum number of descriptors to an item, where each “slot” represents a categorical feature. In this work, we opt to assume the best potential value for each slot is nominal, i.e., mutually exclusive, although there could be more than one suitable value at the same slot for each item (e.g., a song that is both “Rock” and “Blues”). These assigned descriptors also have an inherent order, where a preceding descriptor represents the broader, parent category of its child descriptors. For example, for a item to fit “Glam Rock” descriptor, it’s preceding, broader feature would be “Rock”. As we later show, this formulation allows these features to be natively used as item IDs such as semantic IDs and ID-like ranking features Section 5, as well as regular natural language term-based descriptors in use cases like bag-of-terms retrieval Section 6.1.
5. Main Experiments
In this section, we showcase the performance of the AgenticTagger-produced descriptor features by deploying them as discrete ID features in ML models, in particular, generative recommendation (Section 5.1) and ranking (Section 5.2). Meanwhile, we defer discussion on term-based retrieval (Section 6.1) and interpretability (Section 6.2) to later sections, alongside other analysis-focused research questions.
5.1. Generative Recommendation
5.1.1. Datasets and Evaluation.
We conduct our experiments across three domains from the Amazon Reviews dataset (McAuley et al., 2015): “Sports and Outdoors” (Sports), “Beauty” (Beauty), and “CDs and Vinyl” (CDs), the dataset statistics are shown in Table 6. We adopt the commonly used “last-out” splitting (Kang and McAuley, 2018; Rajput et al., 2023; Hou et al., 2025a), where we hold out the last chronologically interacted item for each user as the test set, and the second-last item as the validation set. Following recent work (Rajput et al., 2023; Hou et al., 2023; Wang et al., 2024a), we evaluate the performance of models with Recall@$K$ and NDCG@$K$, where $K\in\{5,10\}$. To eliminate randomness in model optimization, we repeat our experiments with 3 random seeds and report the average metrics. We use Gemini-2.5-flash and Gemini-2.5-flash-lite for the architect and annotator LLMs, respectively, in our main experiment, and discuss results from Gemini-1.5-flash and Gemini-1.5-flash-lite (specifically, the 001 series) where we achieved similarly competitive performance in Section B.1.
5.1.2. Baselines.
In our main experiment (Table 2), we compare our method with a few streams of models: models with item IDs only, such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019); models enhanced by embedded features: FDSA (Zhang et al., 2019), $\text{S}^{3}$-Rec (Zhou et al., 2020); and models with discrete features (i.e., semantic IDs) derived from semantic information or collaborative information: VQ-Rec (Hou et al., 2023), TIGER (Rajput et al., 2023), HSTU (Zhai et al., 2024), and LLM-Indexer (Jin et al., 2024). We additionally compare our method with models that enhances raw semantic IDs via feature crossing, namely SPM-CID (Singh et al., 2024) and ActionPiece (Hou et al., 2025b) in Figure 3.
5.1.3. Main Results.
We summarize our results on the Amazon Reviews benchmarks in Table 2. Overall, models with sets of discrete item features outperform other competing methods that make use of single-embedding-based item features, while AgenticTagger helps achieve the best overall performance compared to various baselines, demonstrating its effectiveness.
Interestingly, the performance improvements are more significant in the Sports and CDs domain, with 8.9%-11.6% improvements against the closest competing model in NDCG@10, where as the improvement is moderate (2%) in the Beauty domain. We hypothesize that this could be due to AgenticTagger yields the most benefit when it is able to construct its vocabulary from a larger items corpus, where the natural language feedback aggregated via the Distill operation provides more stable signal. We further discuss this hypothesis in Section 7.2, where we show that AgenticTagger benefit from larger dataset with diverse items, even when doing so at the cost of pruning the vocabulary building process.
Figure 3. Comparison against feature-crossing-based methods in NDCG@10. For ActionPiece, we compare against its ablated variant without inference time model ensembling, to isolate the effect on item-content understanding.
5.1.4. Comparisons with ID Crossing.
In light of recent advances on showing the benefit of incorporating dataset collaborative filtering statistics directly into the tokenization process via feature crossing (Hou et al., 2025b; Singh et al., 2024), we also compare our method with two recent tokenization methods that merge frequently co-occurring semantic IDs into new IDs based on observed user behavior sequences, SentencePiece (Singh et al., 2024) and ActionPiece (Hou et al., 2025b). To be specific, these methods represent an alternative path towards better features for recommender systems: automated feature crossing based on user behavior. As shown in Figure 3, AgenticTagger remains a strong tokenization method, demonstrating the effectiveness of leveraging LLMs’ content understanding ability in tokenization in addition to collaborative filtering signals. Meanwhile, we note that AgenticTagger is orthogonal to these crossing-based techniques, and leaves crossing AgenticTagger features to future study.
5.2. Ranking
Table 3. Ranking performance comparison between Item ID baseline and AgenticTagger.
Metric Item ID (Baseline) AgenticTagger (Ours)
R@5 0.0833 $\pm$ 0.0068 0.1205 $\pm$ 0.0080
N@5 0.0589 $\pm$ 0.0051 0.0853 $\pm$ 0.0060
R@10 0.1163 $\pm$ 0.0079 0.1751 $\pm$ 0.0093
N@10 0.0695 $\pm$ 0.0052 0.1028 $\pm$ 0.0061
Besides generative recommendation, another frequent evaluation for recent discrete item ID generation frameworks is ranking (Singh et al., 2024; Zhu et al., 2021). To this end, we evaluate the effect of AgenticTagger-features on a private dataset, since publicly available ranking datasets generally anonymize item features (Zhu et al., 2025b), making AgenticTagger inapplicable. Similar to the generative recommendation experiments Section 5.1, we use the trained ranking model to perform click-through-rate-based ranking, and report performance in Recall@$K$ and NDCG@$K$, where $K\in\{5,10\}$. For dataset construction, we collect interaction logs from a news-feed serving platform over a random subset of 1791 users, 15,830 items, and 47,260 interactions. We report our result as in Table 3, as shown, AgenticTagger brings significant improvement to ranking accuracy compared to raw item IDs, showcasing the effectiveness of our proposed approach in effectively compressing item semantic information.
6. Further Applications
6.1. Term-based Retrieval
Figure 4. Comparison between AgenticTagger (AT) and free-form generated discriptor-tags (FT) variants on term-based retrieval. Top: Recall@k results across baselines. Bottom: Performance-utilization trade-off. AgenticTagger achieves the best cardinality-performance trade-off, with further gain when combined with FT. Our method achieves up to 81.9% feature utilization compared to only 6.1% for free-form generation.
In term-based retrieval, user histories are aggregated by summing the bag-of-terms representation of each interacted item, and target documents are ranked by their dot product scores w.r.t. the user-history representation. We choose this setting since it is a model-free prediction pipeline that naturally handles large cardinalities well, allowing us to study the effect of methods with dramatically varying vocabulary sizes while eliminating randomness in model optimization.
We consider 3 layers of AgenticTagger features (AT), free-form generated descriptor tags (FT), and variants of FT that reduces its vocabulary by grouping semantically similar features via k-means clustering (FT+KM), and finally, combining AT and FT feateures via feature-crossing (AT+FT). For the AT+FT method specifically, we prompt the LLM to generate three free-form keywords in addition to the AgenticTagger vocabulary during the vocabulary assignment stage, to study the potential headroom for combining different methods. The free-form keywords are then crossed with level-2 descriptors in AgenticTagger to form additional features, for effective feature enhancement under moderate vocabulary size overhead. Experiments are conducted on a random subset user interaction logs over a week from the same subset in Section 5.2 that consists of 1791 users, 15,830 items, and 47,260 interactions after removing repeated item consumptions.
The results are shown in Figure 4. By comparing FT+KM and FT, we can see that a larger vocabulary benefits retrieval performance, showing the effectiveness of LLM-generated features. Interestingly, AT can already match the performance of FT, especially on coarse retrieval (R@50), showcasing its advantage in cardinality-sensitive scenarios. Finally, AT+FT achieves the best performance, despite having lower cardinality than vanilla FT (since only three keywords are generated per item), showcasing the potential value of combining AgenticTagger-features with other features such as keywords.
As in prior works for (numerical) discrete item representations for recommendation (Hou et al., 2023, 2025b, 2025a; Singh et al., 2024; Rajput et al., 2023; He et al., 2025), another important property to study for a given vocabulary codebook and its assignment is its vocabulary utilization rate, i.e., how many features are effectively used across the item corpus. While AgenticTagger, by its nature, guarantees that each feature it generates is assigned to at least one item, we believe it is meaningful in this case to examine the number of descriptor features that appear in more than one item, as only these features enable effective cross-item generalization. To this end, we report the vocabulary utilization rate and performance trade-off in the lower sub-figure of Figure 4. Specifically, both AT and AT+FT achieve a better performance-utilization trade-off, demonstrating that AgenticTagger’s design effectively improves the framework beyond free-form generation.
Table 4. Performance improvements after single-turn critique on 5000 samples.
Dataset Metric Vanilla +Critique Improve.
Sports R@5 0.0318 0.0428 +34.59%
N@5 0.0206 0.0293 +42.09%
R@10 0.0500 0.0670 +34.00%
N@10 0.0264 0.0371 +40.31%
Beauty R@5 0.0504 0.0864 +71.43%
N@5 0.0350 0.0602 +72.17%
R@10 0.0738 0.1284 +73.98%
N@10 0.0425 0.0738 +73.65%
CDs R@5 0.0546 0.0666 +21.98%
N@5 0.0374 0.0445 +19.21%
R@10 0.0816 0.0990 +21.32%
N@10 0.0460 0.0549 +19.42%
6.2. Interpretability and Controllability
As discussed earlier, a core advantage of LLM-generated features is their explainability, which naturally suits user-facing use cases, such as situations where one would like to control or constrain the matching mechanism of a recommender system. In this section, we discuss the value AgenticTagger features add to the interpretability and controllability of a recommender system via single-turn critique setting, where a user gives feedback to a recommender system, providing information on the item they would like to interact with next. In particular, this setting is a fundamental ability in many conversational recommender systems (Li et al., 2018; Wang et al., 2022; He et al., 2023; Zhu et al., 2025c), and is commonly manifested in production recommender systems as feedback with keyphrase-based rationales.
Inspired by recent work on using LLMs for user simulation in recommender systems (Bougie and Watanabe, 2025; Park, 2025; Zhang et al., 2025a; Wang et al., 2024e) and conversational recommendation (Wang et al., 2023; Zhu et al., 2025a), we opt to make use of an LLM as user simulator to evaluate whether AgenticTagger can enable a user to perform effective control over recommender system behavior. Specifically, we adopt a communication game with an information bottleneck setting, where an LLM (gemini-2.5-flash in practice) observes the descriptions of target items in its prompt, and is asked to select all potentially relevant top-level (i.e., level 1) features in the vocabulary, which generally consists of a small amount of keywords (e.g., 16 for the CDs and Vinyl domain, the largest public benchmark we evaluate on). We then generate the next item from a pre-trained generative recommender backbone using beam search, following standard practice, except we constrain its search space to be within the allowed level 1 features provided by the user simulator. While this interaction paradigm essentially “leaks” ground-truth information to the recommender, it verifies that (1) AgenticTagger produces a vocabulary to precisely identify meaningful subset of items, and (2) the produced vocabulary aligns well with the semantic understanding of an LLM, which are known to align well with human perception (Chang and Bergen, 2024), demonstrating the interpretability and controllability potential of AgenticTagger features.
As shown in Table 4, this constrained recommendation setting increases model performance significantly across datasets, showcasing that AgenticTagger features effectively communicate the nuances of items in the dataset, and importantly, in an interpretable manner, such that the behavior of the corresponding recommender system using AgenticTagger-features can be straightforwardly controlled by natural-language-bridged critiques.
7. Ablations and Discussions
7.1. Effect of Multi-agent Self-refinement
Figure 5. Distribution of coverage changes between optimization steps across layers. We report level 2-6, which have enough trials to obtain meaningful results for analysis.
In this ablation, we examine the effect of the iterative refinement loop in AgenticTagger on improving item coverages. To verify its effect, we plot the average percentage differences in item coverage (i.e., items that are assigned at least one feature among the current vocabulary) for each clustering job in Figure 5. As shown, the overall trend is that iterative refinement helps improve item coverage, leading to a reduced number of items that cannot be described by the vocabulary as the number of optimization steps grows. While we report the difference between steps 1 and 2 on Sports dataset, the trend is consist across dataset and optimization steps.
On the other hand, we observe that iterative optimization tends to less steadily bring coverage improvements towards later-layer (i.e., finer-grained) AgenticTagger features. To this end, we hypothesize that to effectively improve item coverage, the LLM must be able to understand increasingly nuanced item-item similarities as the hierarchy depth grows; intuitively, it is easier to separate, for example, high-level differences such as Rock music and Blues music and specific subgenres of Rock music.
7.2. Scaling AgenticTagger to Larger Dataset
In AgenticTagger, a key decision to make when scaling to larger datasets is the balance between controlling the maximum branching factor $b$, i.e., how many features an item can associate with at each granularity level, and how much data to subsample for generating the vocabulary using Algorithm 1. While these two techniques are not mutually exclusive, it is nevertheless helpful to understand which tends to yield better performance for a large dataset.
To investigate this trade-off, we compare the quality of AgenticTagger features generated from a random subset of 20k items with a variant of the same algorithm (i.e., Algorithm 1) where we constrain the maximum branching factor $b$ to be 1. That is, we only consider one randomly selected associated feature at each feature level for an item when generating finer-grained features in the next level. We found that surprisingly, constraining the branching factor yields significantly superior performance compared to subsampling, where the performance drops to a comparable level to P5-CID. We hypothesize that constructing the vocabulary from more items helps AgenticTagger assign valid features to long-tail item groups, which are crucial for capturing niche user interests.
7.3. Comparison with Free-form Generation in Generative Recommendation
As we discussed earlier, it is hard to directly make use of free-form generated features for generative recommendation, since the method assigns sets of features to items with no constraints on the number of features per item or the “uniqueness” of the features. To illustrate this point, we discuss an earlier attempt that aims at directly using free-form generated features as semantic ID on Sports domain. We choose this domain since it is a larger domain where model performance is known from prior works.
To be specific, we first prompt an LLM to generate free-form descriptors, and treat any ever-generated descriptors as part of the valid vocabulary. We then train a dual-encoder model that predicts which descriptor is generated for each item among the possible descriptor vocabulary (i.e., every descriptor that has been generated by the LLM), following prior work in LLM-based document topic mining (Xie et al., 2025). This process ensures that the potential match between each descriptor-item pair is sufficiently explored, even when the LLM did not generate a particular descriptor initially. We then perform retrieval-augmented generation with another pass of prompting, except that the model this time is asked to select all relevant descriptors from the top 30 suggested descriptors predicted by the aforementioned dual-encoder model. This process results in a maximum of 27, a minimum of 0, and an average of 12.25 descriptors per item.
Due to the large number of descriptors per item, we then prune the descriptors assigned to each item via two methods. (1) Frequency: we split the tags by their occurrence across the corpus into four equal bins, ignoring tags with frequencies smaller than 10 or larger than 2,000, since these are overly rare or common. Each item’s descriptors are selected by taking the most likely tag from each bin, when available, creating a coarse-to-fine tag sequence similar to semantic IDs. (2) K-means: we select 5,000 centroids by running k-means clustering over the text embeddings of the descriptor vocabulary.
However, this framework resulted in poor overall performance: $NDCG@10$ for frequency-based pruning is 0.185, and for k-means-based pruning it is 0.184, which is unlikely to be comparable to modern semantic ID performance. This demonstrates, in contrast, the effectiveness of AgenticTagger in generating hierarchically structured, semantic-ID-like descriptors.
7.4. Case Study
Table 5. Case Study: AgenticTagger and free-form generation as features. Item content rewritten to preserve privacy.
Item Context Free-form) AgenticTagger
Ground Truth Skipper Limited, L0: Business & Finance
Transformer stock, Order Book Growth, L1: Stock Market
820% 3-yr returns Power Transmission. L2: Individual Performance
User History L0: Business & Finance
(1) Defence bidder, 7k% 5-yr returns (1) Defence PSU, DRDO Bidding. L1: Stock Market
(2) Industrial PSU, 200% 5-mo returns (2) PSU Stocks, Market Returns. L2: Individual Performance, Order Book Analysis
Finally, to understand the improvement brought by AgenticTagger compared to free-form generation, we take an example from our private dataset to analyze the effect of features from both frameworks on connecting the user’s recent interests to the ground-truth item. We analyze a user with two recently interacted items (news feeds, in this case) and the held-out feed content. As shown, the free-form-generated features, while high-quality, failed to generalize across items due to random surface phrase-choices; in contrast, AgenticTagger successfully identified the user’s broad interest in the stock market and detailed interest in stock performance. This performance could be further improved by crossing these features with entity keywords, such as specific stock names and locations.
8. Conclusion
In this work, we introduce AgenticTagger, a general framework for prompting LLMs to generate natural language item-descriptor representations for recommender systems. We showcase the wide applicability of AgenticTagger across various applications, including term-based retrieval, generative retrieval, ranking, and controllability-oriented tasks. We also discuss the properties of AgenticTagger, such as its complexity and feature cardinality. In the future, we plan to investigate methods for improving effectiveness by introducing AgenticTagger realizations that consider collaborative filtering signals during the feature generation phase, following recent success in automated feature crossing.
Appendix A Appendix
A.1. Implementation Details
Table 6. Statistics of public datasets used. “Avg.” denotes the average sequence length.
Datasets #Users #Items #Interactions Avg.
Sports 35,598 18,357 260,739 8.32
Beauty 22,363 12,101 176,139 8.87
CDs 75,258 64,443 1,022,334 14.58
Table 7. Hyperparameter settings for each dataset.
Hyperparameter Sports Beauty CDs
learning_rate 0.005 0.005 0.01
n_semid 6+1 3+1 6+1
vocab_size 2487+80 1103+1145 742+192
warmup_steps 10,000 10,000 10,000
dropout_rate 0.1 0.1 0.1
weight_decay 0.07 0.10 0.07
beam_size 50 50 50
num_layers 4 4 4
d_model 128 128 256
d_ff 1,024 1,024 2,048
num_heads 6 6 6
d_kv 64 64 64
optimizer adamw adamw adamw
lr_scheduler cosine cosine cosine
train_batch_size 256 256 256
max_epochs 200 200 200
early_stop_patience 20 20 20
In this section, we cover our implementation details on public benchmark experiments disclosed in this work.
A.1.1. Baselines and pre-processing.
We directly reuse the data processing implementation open-sourced by Hou et al. 2025b for fair and straightforward comparison. We reuse the reported baseline performance in Hou et al. 2025b, where baseline results for BERTRec, SASRec, FDSA, S³-Rec on Sports and Beauty domain are taken directly from each respective work and rest are re-implementations by Hou et al..
A.1.2. Hyperparameter for GR Training.
We train the generative recommender model from scratch with AgenticTagger as semantic IDs from scratch for a maximum of 200 epochs, and break the training if there are no improvements on $N@10$ for 20 epochs. Following Hou et al. 2025b, we set training batch size to 256 and tune learning rate from $\{1\times 10^{-3},3\times 10^{-3},5\times 10^{-3}\}$ and weight-decay from $\{0.07,0.1,0.15,0.2\}$. We conduct three repeated experiments with random seeds $\{2025,2026,2027\}$, and report the test set performance with the model checkpoints with best average $N@10$ across runs.
A.1.3. GR Inference.
We use standard auto-regressive decoding with a beam size of 20 for inference. However, since AgenticTagger can produce variable length semantic IDs, we do not sample from the base generative recommender for a fixed number of step, and instead generate until we reach the end of sequence token. The generated sequence are then ranked by their beam scores.
A.1.4. AgenticTagger
By default, we choose 15 as the initial cluster size across all datasets and number of refinement cycles to 3. The minimum cluster size for sub-clustering is set to 30. We break the iterative refinement loop if the item coverage for any cluster is greater than 95, or the refiner LLM fail to generate more than 20 tickets. Due to the large number of items for CDs domain, we set the sub-clustering threshold to 100 for branching-controlled version of the algorithm, and 500 for the subsample variant. While the final vocabulary size of AgenticTagger due to stochasticity of program execution, we report the base vocabulary and the number of conflict resolvers (i.e., resolving token in case two items hashed with the same set of AgenticTagger tags) as “base tokens + resolvers” in Table 7 as “vocab_size.” As shown, AgenticTagger is significantly more efficient than its semantic ID counter-parts using less than 3000 words across the board on academia benchmarks.
Appendix B Further Related Works
ML Models with Natural Language Parameters. Beyond recommender systems, there has been continuing effort for introducing parameters tied to explicit natural language concepts into ML models, e.g., concept-bottleneck models (Koh et al., 2020; Sun et al., 2025), typically for its improved interpretability (Zhong et al., 2024). However, this stream of work generally don’t aim for improved accuracy, and frequently assumes there exists a set of concepts (in our case, features) that are readily available (Koh et al., 2020). The most relevant work to ours is perhaps label-free concept bottleneck model (Sun et al., 2025), which prompts a LLM with task description for concepts. However, our aim is to infer a set of features from a large sets items for recommendation, the nature of which is typically hard to infer from the task description itself, unlike domains such as image classification with fixed set of labels.
LLM for Text Mining At Scale Another relevant stream of work to AgenticTagger, as discussed through the paper, is recent advance in leveraging LLMs for scalable text mining, typically minifest as topic/latent theme mining (Xie et al., 2025; Pham et al., 2024) or taxonomy mining (Wan et al., 2024). However, these framework generally do not consider downstream applications, and usually imposes sequential dependencies on the interaction between LLMs and text corpus (Wan et al., 2024; Pham et al., 2024), where as AgenticTagger extensively use parallelize operation, making scaling to recommendation-scale item corpus feasible.
B.1. Effect of Base Models
In our main experiments on public benchmarks, we opt to use the gemini-2.5 family of models since they are released closest to the time the experiments were conducted. However, to investigate if our framework are robust to the choice of based models, we also experimented with gemini-1.5-flash and gemini-1.5-pro as annotator and architect LLM, respectively. As shown in Table 8, both models yields competitive performance, while gemini-2.5 maintains a slight advantage across datasets, likely due to its stronger overall abilities in text generation and understanding.
Table 8. Impact of the backbone model on performance. Gemini-2.5 yields slightly better performance across domains, while Gemini-1.5 yields comparable results.
Backbone LLM Sports Beauty CDs
R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10
Gemini-1.5 0.0284 0.0184 0.0448 0.0237 0.0469 0.0319 0.0710 0.0397 0.0531 0.0349 0.0812 0.0439
Gemini-2.5 0.0299 0.0194 0.0474 0.0251 0.0492 0.0332 0.0728 0.0408 0.0544 0.0358 0.0836 0.0453
B.2. Sample Prompts (Sports Domain)
Architect - Generate Inital Vocabulary You are an expert taxonomy architect. Your task is to create a set of fine-grained sub-categories.
The items you are categorizing all belong to the parent category: "{parent_rule_description}".
Your new categories MUST be more specific subdivisions of this parent category. Do not simply repeat the parent category’s name. For example, if the parent is "Firearm Accessories", your sub-categories should be "Optics", "Holsters", "Cleaning Kits", etc., not "Firearm Accessories" again.
Primary Goal: Create a set of ˜{n_target_rules} mutually exclusive sub-categories that logically partition the parent.
{context_prompt}
Input Data:
Below are diverse product examples. Use them to understand the product landscape you need to partition.
{sample_text}
Task & Strict Output Format:
Devise a set of categories. For each, provide a "name" and a "description" with "INCLUDES" and "EXCLUDES" clauses. Respond ONLY with a single JSON object: {"categories": [{...}]}.
Your JSON Response:
Annotator {context_prompt}
Above, I have provided the parent category and some product examples.
Your Goal:
Propose a list of exactly
{n_target_rules} sub-categories.
Constraint - Overlap:
Sub-categories must be mutually exclusive.
Constraint - Coverage:
Together, they should cover most common items found in "
{parent_rule_description}".
Constraint - Specificity:
Do NOT create a "Miscellaneous" or "Other" category. Every category must have a specific, meaningful name.
Constraint - Style:
Use professional, industry-standard terminology.
Output Format:
Your response must be a JSON object with the following structure:
"categories": [
"name": "Category Name",
"description": "Definition of what is included",
"includes": ["Example 1", "Example 2"],
"excludes": ["Non-example 1"]
Annotator - Generating Error Feedbacks You are a data analyst specializing in taxonomy quality control. Analyze a cluster of {len(ticket_cluster)} uncategorized products and synthesize a single, structured change proposal.
Existing Category Rules:
{existing_rules_text}
Analysis of Uncategorized Items:
{ticket_examples_text}
Task: Decide the most logical action and formulate a proposal. Respond ONLY with a single JSON object. Your response MUST strictly follow one of the formats below.
1. To create a new category:
{"change_type": "CREATE_NEW_CATEGORY", "problem_summary": "<summary>", "suggested_change": {"new_rule_description": "New Category Name: INCLUDES: ... EXCLUDES: ..."}}
2. To expand an existing category (THIS IS A VERY STRICT FORMAT):
The "refined_description" MUST be a single complete string, starting with a name, then "INCLUDES", then "EXCLUDES".
GOOD EXAMPLE:
"change_type": "EXPAND_EXISTING_CATEGORY",
"problem_summary": "The current ’Outdoor Gear’ rule is too generic and should explicitly include tactical accessories.",
"suggested_change": {
"rule_id_to_refine": "rule_a4368cef",
"refined_description": "Outdoor & Tactical Gear: INCLUDES: Tents, backpacks, sleeping bags, and tactical accessories like gloves, belts, and pouches. EXCLUDES: Specialized sporting equipment, firearms, and knives."
3. To ignore outliers:
{"change_type": "IGNORE_AS_OUTLIERS", "problem_summary": "<summary>", "suggested_change": {"reason": "<reason>"}}
Your JSON Response:
Architect: Taxonomy Refinement You are a senior taxonomy manager. Review the following change proposals and decide whether to approve or reject each. Your goal is a clean, non-overlapping taxonomy. But if the proposal is reasonable, you should try to accomodate that. You should also reject proposals that are out-of-place by common sense. Respond ONLY with a JSON list of objects, each with "proposal_id", "decision" (’APPROVED’ or ’REJECTED’), and "reasoning".
Proposals for Review:
{proposals_text}
Your JSON Decision List:
B.3. Tags Generated
Table 9. Discovered Vocabulary (Sub-sampled) Aross Domains
Beauty Domain Music (CDs) Domain Sports Domain
L1: Skincare - Facial Treatments L1: Jazz Music L1: Fitness & Exercise Equipment
L1: Haircare - Treatments & Styling L1: Electronic Dance Music (EDM) L1: Camping & Hiking Gear
   L2: Hair & Scalp Treatments    L2: Smooth Jazz    L2: Shelter Systems
   L2: Hair Styling Tools & Acc.    L2: Free Jazz & Avant-Garde    L2: Sleeping Systems
      L3: Hair Dryers       L3: Interdisciplinary Avant-Garde       L3: Sleeping Bag Liners
      L3: Flat Irons & Straighteners       L3: Free Improv. & Avant-Garde       L3: Sleeping Bags
         L4: Travel & Mini Flat Irons          L4: Free Jazz (Post-Bebop)          L4: Sleeping Bag Shapes
         L4: Professional Salon Irons          L4: Avant-Garde (Exp.)          L4: Temp.-Rated Bags
            L5: Specialty Function             L5: Atonal Exploration             L5: Mild to Cool Bags
            L5: Ceramic Plate Irons             L5: Genre-Infused Exp.             L5: Cold Weather Bags
               L6: Adv. Plate Tech.                L6: 10-30 Degree Bags
               L6: Ionic Technology                L6: 0-Degree Rated Bags
We show qualitative examples of the discovered vocabulary as in Table 9. For presentation purpose, we pick two descriptors per level, but only expand on one of them, until no valid children was found. Note that in AgenticTagger, the depth of each hierarchy path could be variable length, such as we show in the CDs domain example. As showcased in the table, AgenticTagger are able to find clear separations of products, especially towards the earlier layers of the hierarchy (e.g., “Skincare” and “Haircare” products). However, towards the end of the hierarchy, the boundary between products becomes fuzzy, and it is more challenging to come up with clean, disentangled separation rules (e.g., “Specialty Function” and “Ceramic Plate Irons”), and it becomes harder to enforce a nominal constraint on features. While we partially address this issue by instruction the LLM to select best fitting descriptor for each product when multiple are technically correct, developing method to fundamentally address such disentanglement issue remains an open problem. Nevertheless, even at later layers, AgenticTagger could still discover meaning descriptors that separates the products, such as sleeping bags by their applicable temperature range as shown in the Sports domain.
B.4. Effect of Multi-agent Self-refinement on Other Domains
Figure 6. Coverage improvements across domains. We discuss further details in Section B.4.
We showcase the improvements of coverage by reporting the distribution of delta changes in coverage rate (i.e., the number of items that can be described by at least one descriptor) for each clustering job $Q$ in Figure 6. The trend is consistent across domains, where (1) our designed multi-agent self-refinement loop effectively refines the vocabulary, improving item coverage; and (2) the trend of improvement (i.e., delta greater than zero) becomes less stable in later layer, possibility due to it is more challenging to generate disentangled descriptors, as we qualitatively analyzed in Section B.3. Regardless, wilcoxon signed-rank test show improvements are significant (i.e., delta greater than zero) with p-value smaller than 0.05, except for level five and six on CDs domain, corroborating the general effectiveness of the self-refinement mechanism.