### v2

# AgenticTagger: Generating Structured Item Representation for Recommendation with LLM Agents

Abstract.
High-quality representations are a core requirement for effective recommendation. In this work, we study the problem of LLM-based descriptor generation, i.e., keyphrase-like natural language item representation generation frameworks with minimal constraints on downstream applications. We propose AgenticTagger, a framework that queries LLMs for representing items with sequences of text descriptors. However, open-ended generation provides little control over the generation space, leading to high cardinality, low-performance descriptors that render downstream modeling challenging. To this end, AgenticTagger features two core stages: (1) a vocabulary-building stage in which a set of hierarchical, low-cardinality, and high-quality descriptors is identified, and (2) a vocabulary-assignment stage in which LLMs assign in-vocabulary descriptors to items. To effectively and efficiently ground vocabulary in the item corpus of interest, we design a multi-agent reflection mechanism in which an architect LLM iteratively refines the vocabulary guided by parallelized feedback from annotator LLMs that validate the vocabulary against item data. Experiments on public and private data show AgenticTagger brings consistent improvements across diverse recommendation scenarios, including generative and term-based retrieval, ranking, and controllability-oriented, critique-based recommendation.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
Figure 1. Illustration of the proposed method. Items are represented using an ordered sequence of LLM-generated descriptors that compress task-relevant information.
1. Introduction
High-quality item representations are the foundation of effective recommendation systems (Rendle et al., 2009; Kang et al., 2021; Rajput et al., 2023; Singh et al., 2024; Hou et al., 2023; Coleman et al., 2023; Zhang et al., 2024b, 2025b). These representations are commonly derived from collaborative filtering (Rendle et al., 2009; Koren et al., 2009; Kabbur et al., 2013a; He et al., 2017; Kang and McAuley, 2018) and/or content signals (Sharma et al., 2015; Mooney and Roy, 2000; Rajput et al., 2023; Singh et al., 2024; Hou et al., 2023, 2025b, 2025a; Wang et al., 2024e) via representation learning. Recently, there have been numerous attempts to enhance item representations with LLMs, typically by generating auxiliary unstructured item descriptions (Valizadeh et al., 2025; Ren et al., 2024; Shi et al., 2025; Yada and Yamana, 2024; Wang et al., 2024c) or prompting LLMs to infer user interests from interacted items and directly make recommendations (Liu et al., 2025a, b; Jalan et al., 2024; Jiang et al., 2025; Wang et al., 2024d). However, these approaches either require modeling high-cardinality unstructured natural language descriptions or deploying specialized recommendation frameworks, whereas large-scale recommender systems frequently benefit from low-cardinality, broadly applicable item representations such as hashed categorical features (Kang et al., 2021; Coleman et al., 2023; Singh et al., 2024; He et al., 2025; Hou et al., 2023, 2025b, 2025a). Thus, despite recent advances in both LLM-enhanced and structured item representations, it is unclear how to effectively leverage LLMs’ strong semantic understanding ability to enhance structured item representations.
Figure 2. The AgenticTagger Framework. AgenticTagger exploits LLMs’ content-understanding ability to automatically construct an interpretable, hierarchical descriptor vocabulary from data. Then, relevant descriptors in the vocabulary are assigned to corresponding items by querying LLMs in parallel over the item corpus. These features could then power various downstream applications, such as generative retrieval, ranking, and critique-based recommendation.
To this end, we study item-side, LLM-based generation frameworks for discrete feature representations (i.e., tag-like natural language descriptors), capable of exploiting LLMs’ reasoning ability to produce compact, structured, and fixed-slot nominal features with minimal constraints on downstream models, as shown in Figure 1. We argue that this setting holds three complementary values compared to its alternatives, such as generating unstructured item descriptions with LLMs or prompting LLMs directly for recommendations. First, discrete features such as hashed item IDs and text descriptors provide irreplaceable value in recommender systems, due to their compatibility with diverse architectures (Zhang et al., 2024b, 2025b; Wang et al., 2017, 2021; Kabbur et al., 2013b; Cheng et al., 2016; Song et al., 2019), their effectiveness in improving recommendation accuracy (Rajput et al., 2023; Hou et al., 2023, 2025b; Singh et al., 2024), and their scalability when deployed in keyword-based retrieval or advertisement targeting (Wartena et al., 2010; Google Ads Help, 2026). Second, item representation scales easily in the typical case where a recommender system serves more users than items, and the item corpus remains more stable than user interests (McAuley et al., 2015; Gao et al., 2022). Finally, as we discuss in a later section (Section 5.3), text-based descriptor representations provide inherent interpretability and are well-suited for user-facing applications.
While existing efforts typically directly generate unstructured item descriptions from LLMs (Valizadeh et al., 2025; Ren et al., 2024; Shi et al., 2025; Yada and Yamana, 2024; Wang et al., 2024c), free-form LLM generations offer no control over their exact word choices, and thus there is no way to enforce structural constraints. For example, it is challenging to enforce mutual exclusivity between descriptions from free-form generation (Section 6.3), making these descriptors unsuitable for our use case. Further, a good set of text descriptors should ideally capture meaningful frequent patterns in the item corpus (Xie et al., 2025; Rajput et al., 2023; Singh et al., 2024). Yet, recommender system corpora generally exceed the context limit of LLMs (McAuley et al., 2015; Harper and Konstan, 2015), and thus it is challenging to incorporate the abstract constraint of generating descriptors that meaningfully highlight item similarities and differences w.r.t. other items, as this requires a global view of the entire item corpus. Finally, as LLMs can pick any arbitrary surface form (e.g., ”Blues music” and ”Blues great hits”) for the same underlying concept in free-form generation, such an approach often leads to vocabulary explosion, where most generated descriptors occur only once across the item corpus, as we discuss in Section B.3. This renders many features unusable without further processing, compromising downstream recommendation performance (Section 6.3).
To address these issues, we propose AgenticTagger, a multi-LLM framework that queries LLMs for natural language descriptors (i.e., tag-like descriptors shown in Figure 1), where the descriptors come from an automatically learned low-cardinality, hierarchically structured vocabulary. Specifically, AgenticTagger features two key stages: (1) vocabulary building (Section 4.1), where AgenticTagger automatically learns a hierarchical coarse-to-fine descriptor taxonomy, drawing insights from LLMs for taxonomy mining (Wan et al., 2024; Xie et al., 2025; Pham et al., 2024) and codebook learning for item representation in recommendation (Hou et al., 2023, 2025b, 2025a; Singh et al., 2024; He et al., 2025; Rajput et al., 2023); and (2) vocabulary assignment (Section 4.2), which prompts LLMs to assign relevant descriptors from the vocabulary to items, constraining their generation space. By mining a high-quality, low-cardinality descriptor vocabulary and performing in-vocabulary generation, AgenticTagger effectively constrains the LLMs’ generation space while equipping LLMs with a global, hierarchical view of the item corpus, enforcing structural constraints on LLM-generated descriptors.
Under the general AgenticTagger framework, a core design challenge is how to effectively “propagate” signals through LLM generations during the vocabulary-building stage to improve the quality of the vocabulary and better cover diverse items in the item corpus. To this end, following the numerous recent successes in self-improving LLM agents through verbal feedback (Shinn et al., 2023; Majumder et al., 2024; Xie et al., 2024), we adopt an LLM-based refinement loop, which refines the vocabulary by aiming for better coverage of the item corpus through LLM-generated failure reports on items that do not fit into the current vocabulary. Unlike these prior works, AgenticTagger uniquely adopts a multi-agent architect-annotator framework, where a stronger (yet slower) architect LLM maintains the descriptor vocabulary, while numerous copies of the weaker (yet faster) annotator LLM scan through the item corpus in parallel, making AgenticTagger scalable.
We conduct experiments on AgenticTagger across public benchmarks and a private dataset, showing the broad applicability of AgenticTagger via its improvements to diverse recommendation settings, including generative recommendation and ranking. We also show that these natural language descriptors natively bring interpretability and controllability to recommender systems, such as by making generative recommender models compatible with critiques by controlling their decoding process via user feedback. Our contributions are as follows:
  • To the best of our knowledge, we are the first to apply agentic frameworks to feature mining to benefit recommender systems, opening up a new direction for future work.
  • We identify a core challenge in LLM-based structured item representation generation, namely, the lack of vocabulary control, and propose a multi-agent, two-stage descriptor generation framework to address this issue.
  • We show the value of LLM-agent-produced, structured natural language item representations in recommender systems, for their performance, interpretability, and flexibility.
2. Related Work
LLM-generated Features in RecSys. Prior work in introducing LLM-produced features in recommender systems broadly falls into two streams: LLM-as-encoders (Zhang et al., 2024b, 2025b) and LLM-as-text-generators (Ren et al., 2024; Yada and Yamana, 2024; Valizadeh et al., 2025; Liu et al., 2024, 2025a), and this work aims at enhancing the latter. In contrast to the popular goal of generating user profiles that improve an LLM itself at the recommendation task (Liu et al., 2025a, b), we consider specifically the case where the recommendation model is a general feature-aware ML model, which aligns well with most existing deployed recommendation models. Compared to the most similar works that encode LLM-produced descriptions (Ren et al., 2024; Yada and Yamana, 2024; Valizadeh et al., 2025), we are the first to study fixed-slot, nominal feature generation, and the first to demonstrate the value of LLM agents in feature generation for RecSys.
Discrete Representation in RecSys. Our work is also similar to a series of efforts in building better discrete features for RecSys, which typically involve discretizing item representations with product quantization and its variants (Jegou et al., 2011; Ge et al., 2014; Hou et al., 2023, 2025a), RQ-VAE (van den Oord et al., 2017; Rajput et al., 2023), and K-means (Ju et al., 2025; Zhang et al., 2024c). However, these methods do not leverage the reasoning ability of LLMs. As we will later show in the paper, discrete features produced from raw text by exploiting LLMs’ content understanding ability outperform their counterparts, such as RQ-VAE-based semantic IDs.
LLM Agents for RecSys. There have been numerous recent efforts in leveraging LLM agents to improve RecSys. These frameworks generally focus on leveraging LLMs for item recommendation (Shu et al., 2023; Zhao et al., 2024; Thakkar and Yadav, 2024; Wang et al., 2024b; Xia et al., 2026; Huang et al., 2025), or enabling better interactivity by incorporating LLMs and LLM agents into the recommendation-serving stage (Friedman et al., 2023; He et al., 2023; Zhang et al., 2024a). However, these frameworks typically require specialized pipelines, whereas AgenticTagger focuses on feature generation, which potentially benefits diverse downstream recommendation models. Meanwhile, to the best of our knowledge, AgenticTagger is the first framework to study agentic feature generation for recommender systems.
3. Problem Statement
Consider the standard recommendation setting with item features. Given a set of users $\mathcal{U}$ and items $\mathcal{I}$, the objective is to predict unobserved items of interest for a user $u\in\mathcal{U}$ based on their history of interactions $S_{u}\subset\mathcal{I}$. Each item $i\in\mathcal{I}$ is associated with raw textual attributes $X_{i}\in\mathcal{X}$, where $\mathcal{X}$ is the original feature space. In the context of feature generation, our goal is to find an optimal one-to-many transformation function, $g:\mathcal{X}\to\mathcal{P}(A)$, that maps these original attributes to a new discrete feature space $A$. The optimal transformation, $g^{*}$, is one that maximizes the performance of a downstream recommender system, $f$:
$g^{*}=\arg\max_{g}\mathcal{M}(f_{\theta^{*}(g)})$
where $f_{\theta^{*}(g)}$ is the recommender model trained using the generated features, and $\mathcal{M}$ represents relevant performance metrics, e.g., offline metrics and key business metrics. Note that while AgenticTagger does not directly optimize for end performance due to the challenge of propagating gradients through an API-supported LLM, the best vocabulary should ideally, in our problem formulation, maximize downstream task performance. We provide an overview of the notations used in Table 10.
4. AgenticTagger
As shown in Figure 2, the overall pipeline of AgenticTagger involves two key stages: the vocabulary-building stage, where we produce a descriptor vocabulary by observing typical items in the dataset, and the vocabulary-assignment stage, where we assign the best set of descriptors to each item.
4.1. Vocabulary Building
Algorithm 1 Vocabulary Building
Input: Item corpus $\mathcal{I}$, LLM Ensemble $\Pi$
Hyperparameters: Max hierarchy depth $D_{max}$, split threshold $\tau_{split}$
Output: Hierarchical descriptor set $A_{all}$
$A_{all}\leftarrow\emptyset$
$p_{0}\leftarrow\text{new Descriptor}(\text{items}\leftarrow\mathcal{I},\text{parent}\leftarrow\text{null})$
$Q\leftarrow\{p_{0}\}$
for each $d\in\{1,\dots,D_{max}\}$ do
  if $|Q|=0$ then
   break
  end if
  $Q_{next}\leftarrow\emptyset$
  for each parent feature $p\in Q$ do
   if $|p.\text{items}|\geq\tau_{split}$ then
[…]
    $A_{all}\leftarrow A_{all}\cup A_{children}$
    $Q_{next}\leftarrow Q_{next}\cup A_{children}$
   end if
  end for
  $Q\leftarrow Q_{next}$
end for
return $A_{all}$
During the vocabulary-building stage, AgenticTagger automatically learns a vocabulary from the item corpus. However, gradient-based vocabulary learning (Rajput et al., 2023; Hou et al., 2023, 2025a; Zhang et al., 2024c) is inapplicable to prompting-based LLM agents, and thus the common solution is prompting an LLM in mini-batches (Wan et al., 2024; Pham et al., 2024) and refining an in-context vocabulary via LLM self-refinement (Madaan et al., 2023). Yet, such an iterative approach involves a non-parallelizable loop over the whole item corpus, as LLMs must sequentially refine their in-context vocabulary after seeing each batch of items, which scales poorly to large datasets (Xie et al., 2025).
4.1.1. Hierarchical Vocabulary Construction.
In light of the above issues, in AgenticTagger, we opt to scale existing LLM-based frameworks by making the vocabulary-building process parallelizable, where multiple LLMs are deployed concurrently to update a shared vocabulary that describes observed items. An overview of this process is illustrated in Algorithm 1 and Algorithm 2. Specifically, the iterative feature generation follows a hierarchical-clustering-like pipeline (Algorithm 1), where we generate an initial set of features that partitions the whole item corpus, followed by a top-down procedure to generate finer-grained features that further partition items associated with each feature into smaller subsets. For example, if the initial features include high-level category information such as “CDs and Vinyl”, all relevant items will then form a subcluster, for generating finer-grained features such as “Rock” and “Blues”, as shown in Figure 2-(i). By decomposing the task of generating an entire vocabulary into a top-down hierarchical procedure, AgenticTagger effectively limits the number of descriptors an LLM needs to produce at a time, which scales better to a large corpus. Meanwhile, the resulting hierarchical vocabulary naturally organizes items in a coarse-to-fine taxonomy, similar to hierarchical semantic IDs (Rajput et al., 2023; He et al., 2025), which are known to act as effective item representations.
4.1.2. Multi-agent Self-refinement.
At the core of this top-down generation loop is a multi-agent iterative refinement (Figure 2-(ii) and Algorithm 2) subroutine reused for each feature generation job $Q_{i}$, which aims to produce a set of features that describe common items in the corresponding subset. Our core insight is leveraging a stronger (though potentially slower) LLM, $\pi_{Architect}$, to generate and refine the vocabulary subset relevant to each $Q_{i}$, and copies of a faster (and potentially weaker) LLM, $\pi_{Annotator}$, which attempt to annotate the current item corpus using the available vocabulary and provide feedback to $\pi_{Architect}$ to improve the vocabulary. Unlike in single-LLM self-refinement (Wan et al., 2024; Pham et al., 2024; Madaan et al., 2023), the most costly $\pi_{Annotator}$ operations are parallelized in AgenticTagger, while we can still use their feedback to prompt $\pi_{Architect}$ to refine the vocabulary.
As shown in Algorithm 2, during the subroutine execution, we first prompt the $\pi_{\text{Architect}}$ LLM to generate the initial set of descriptors by observing a small subset of items, $\text{Distill}(I_{sub})$, in its context, and then deploy copies of (or batched queries to) the annotator LLMs $\pi_{\text{Annotator}}$ to determine if any of the currently known features apply to each item in $I_{sub}$, generating a set of error suggestions $\mathcal{E}$. The specific prompts are as shown in Section B.1. The error reports $\mathcal{E}$ then act as inductive biases for prompting the vocabulary maintainer LLM $\pi_{\text{Architect}}$ to refine the current set of features. To avoid overcrowding $\pi_{\text{Architect}}$’s context, $\text{Distill}(\cdot)$ refers to a general operation that selects diverse representative samples from a larger set of item descriptions, such as the centroids from K-Medoids clustering (Kaufman and Rousseeuw, 1987), as we use in our experiments. We repeat this iterative refinement loop until a fixed number of optimization steps has been reached, or a desired number of items in $I_{sub}$ can be described by the generated features $A$. The union of features generated for each subset of the item corpus then forms the final feature space $A_{all}$, providing a coarse-to-fine description for the whole item corpus.
Algorithm 2 Multi-agent Self-refinement
Input: Item corpus subset $I_{sub}\subseteq\mathcal{I}$;     LLMs $\Pi=\{\pi_{\text{Architect}},\pi_{\text{Annotator}}\}$
Hyperparameters: Max refinement cycles $C_{max}$, anomaly threshold $\tau_{anom}$
Output: Descriptor vocabulary AA
$A\leftarrow\pi_{\text{Architect}}(\text{Distill}(I_{sub}))$
for $c=1$ to $C_{max}$ do
  $(I_{assigned},I_{unassigned},\mathcal{E})\leftarrow\text{ParallelAssign}(I_{sub},A,\pi_{\text{Annotator}})$
  if $|I_{unassigned}|<\tau_{anom}$ then
   break
  end if
  $A\leftarrow\pi_{\text{Architect}}(A,\text{Distill}(\mathcal{E}))$
end for
return AA
4.2. Vocabulary Assignment
Once the feature vocabulary $A_{all}$ is built, the vocabulary assignment stage for AgenticTagger is straightforward, where we simply constrain the descriptors output by the LLMs to fall into the vocabulary by prompting, i.e., $g(X_{i})=\{a\in A\mid\pi(X_{i},A)\to a\}$ for any raw item attribute $X_{i}$. For example, in AgenticTagger, for the sake of implementation simplicity, we opt to reuse the same operation from building the vocabulary, $\text{ParallelAssign}(I,A_{all},\pi_{\text{Annotator}})$, where an LLM is prompted to assign features to an item by observing all features in its context, exploiting modern LLMs’ long-context and reasoning abilities (Zhao et al., 2025). As we later show in Section 5, the generated descriptors are indeed of sufficient quality to support good performance in downstream recommendation tasks. We provide further analysis and qualitative examples in Section B.4.
We note that in principle, there might be multiple descriptors of fit at each depth of the vocabulary hierarchy (e.g., a CD that is both “Rock” and “Blues”). To this end, we simply instruct the LLM to select a single best descriptor per level, enforcing structural constraints while preserving item information. As we later showcase in our experiments (Section 5), this is sufficient for good downstream recommendation performance, and we leave better collision-resolving methods to future work.
4.3. Complexity Analysis
4.3.1. Vocabulary Generation.
In AgenticTagger’s automated vocabulary-building algorithm, the parallel calls to the annotator LLMs during the vocabulary creation phase dominate. In its most basic version, the algorithm will incur $O(C_{max}\cdot N\cdot b^{D_{max}})$ LLM calls, assuming an item will be assigned to an average of $b$ features at each depth level $d\in\{1,\dots,D_{max}\}$, which grows quickly as the depth of the feature hierarchy grows. However, the expressiveness of the feature hierarchy also grows exponentially w.r.t. the depth of the hierarchy, similar to other hierarchical discrete item representation methods (Rajput et al., 2023; Zhang et al., 2024c), and thus a few layers are often sufficient; as we later show in our experiments, 3-6 layers of features are sufficient for good performance across public benchmarks (Table 1).
Further, since the goal of vocabulary building is solely to generate a set of feature descriptors, we can also heuristically constrain the branching factor by limiting how many features an item is assigned to at each layer, such as randomly assigning an item to one of its associated features during the branching operation. In practice, we observe that a branching factor of 1 is also sufficient for good performance (Section 6.2), in which case the complexity w.r.t. LLM calls becomes $O(C_{max}\cdot N\cdot D_{max})$. Finally, compared to prior LLM-based document categorization methods (Wan et al., 2024; Pham et al., 2024), since AgenticTagger does not have sequential dependency between most of its dominant LLM operations (specifically, the $\pi_{\text{Annotator}}$ annotations), the overall execution time can be further reduced approximately linearly w.r.t. the number of concurrent threads in the ParallelAssign operation.
Table 1. AgenticTagger for Generative Recommendation on academic benchmarks. For each metric, the best result is in bold and the second-best is underlined. Our model’s performance is reported across 3 runs. We provide additional comparison against crossing-based IDs and free-form generations in Figure 3 and Section 6.3.
Datasets Metric ID-based Feature + ID Generative
BERT4Rec SASRec FDSA S³-Rec VQ-Rec LMIndexer HSTU P5-CID TIGER AgenticTagger
Sports R@5 0.0115 0.0233 0.0182 0.0251 0.0181 0.0222 0.0258 0.0287 0.0264 0.0299 $\pm$0.0010
N@5 0.0075 0.0154 0.0122 0.0161 0.0132 0.0142 0.0165 0.0179 0.0181 0.0194 $\pm$0.0007
R@10 0.0191 0.0350 0.0288 0.0385 0.0251 0.0414 0.0426 0.0400 0.0474 $\pm$0.0002
N@10 0.0099 0.0192 0.0156 0.0204 0.0154 0.0215 0.0224 0.0225 0.0251 $\pm$0.0004
Beauty R@5 0.0203 0.0387 0.0267 0.0387 0.0434 0.0415 0.0469 0.0468 0.0454 0.0492 $\pm$0.0007
N@5 0.0124 0.0249 0.0163 0.0244 0.0311 0.0262 0.0314 0.0315 0.0321 0.0332 $\pm$0.0008
R@10 0.0347 0.0605 0.0407 0.0647 0.0741 0.0704 0.0701 0.0648 0.0728 $\pm$0.0010
N@10 0.0170 0.0318 0.0208 0.0327 0.0372 0.0389 0.0400 0.0384 0.0408 $\pm$0.0008
CDs R@5 0.0326 0.0351 0.0226 0.0213 0.0314 0.0417 0.0505 0.0492 0.0544 $\pm$0.0005
N@5 0.0201 0.0177 0.0137 0.0130 0.0209 0.0275 0.0326 0.0329 0.0358 $\pm$0.0002
R@10 0.0547 0.0619 0.0378 0.0375 0.0485 0.0638 0.0785 0.0748 0.0836 $\pm$0.0003
N@10 0.0271 0.0263 0.0186 0.0182 0.0264 0.0346 0.0416 0.0411 0.0453 $\pm$0.0002
4.3.2. Vocabulary Assignment.
Once the vocabulary is built, the assignment stage simply involves looping through all items of interest, and the assignment loop induces $O(|\mathcal{I}|)$ LLM calls, i.e., the cost of querying $\pi_{Annotator}$ in a single-pass manner over the item corpus. To this end, after the vocabulary is built, AgenticTagger can easily scale to large item corpora with parallelized calls to LLMs. Further, since the only difference across LLM calls during the assignment stage is the item description, the inference can, in principle, be further accelerated by infrastructure optimizations such as KV-Cache (Pope et al., 2023), which we leave for future work.
4.4. Modeling AgenticTagger Features
Once the process is complete, AgenticTagger assigns a fixed maximum number of descriptors to an item, where each “slot” represents a categorical feature. In this work, we opt to assume the best potential value for each slot is nominal, i.e., mutually exclusive, although there could be more than one suitable value at the same slot for each item (e.g., a song that is both “Rock” and “Blues”). These assigned descriptors also have an inherent order, where a preceding descriptor represents the broader, parent category of its child descriptors. For example, for an item to fit the “Glam Rock” descriptor, its preceding, broader feature would be “Rock”. As we later show, this formulation allows these features to be natively used as item IDs, such as semantic IDs and ID-like ranking features (Section 5).
5. Main Experiments
In this section, we showcase the performance of the AgenticTagger-produced descriptor features by deploying them as discrete ID features in ML models, in particular, generative recommendation (Section 5.1) and ranking (Section 5.2). We then discuss an interpretability-oriented application, namely critique-based generative recommendation (Section 5.3), before other ablation studies and analyses.
5.1. Generative Recommendation
5.1.1. Datasets and Evaluation.
We conduct our experiments across three domains from the Amazon Reviews dataset (McAuley et al., 2015): “Sports and Outdoor” (Sports), “Beauty” (Beauty), and “CDs and Vinyl” (CDs); the dataset statistics are shown in Table 6. For the larger CDs domain only, we constrain AgenticTagger’s branching factor to 1 for scalability and apply the vanilla AgenticTagger algorithm to the other two domains (see Section 6.2 for discussions). We adopt the commonly used “last-out” splitting (Kang and McAuley, 2018; Rajput et al., 2023; Hou et al., 2025a), where we hold out the last chronologically interacted item for each user as the test set, and the second-last item as the validation set. Following recent work (Rajput et al., 2023; Hou et al., 2023; Wang et al., 2024a), we evaluate the performance of models with Recall@$K$ and NDCG@$K$, where $K\in\{5,10\}$. To eliminate randomness in model optimization, we repeat our experiments with 3 random seeds and report the average metrics. We use Gemini-2.5-flash and Gemini-2.5-flash-lite for the architect and annotator LLMs, respectively, in our main experiment. We discuss results from Gemini-1.5-flash and Gemini-1.5-flash-lite (specifically, the 001 series), which achieved similarly competitive performance, in Section 6.4.
5.1.2. Baselines.
In our main experiment (Table 1), we compare our method with several families of models: models with item IDs only, such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019); models enhanced by embedded features: FDSA (Zhang et al., 2019), $\text{S}^{3}$-Rec (Zhou et al., 2020); and models with discrete features (i.e., semantic IDs) derived from semantic information or collaborative information: VQ-Rec (Hou et al., 2023), TIGER (Rajput et al., 2023), HSTU (Zhai et al., 2024), and LLM-Indexer (Jin et al., 2024). We additionally compare our method with models that enhance raw semantic IDs via feature crossing, namely SPM-CID (Singh et al., 2024) and ActionPiece (Hou et al., 2025b) in Figure 3.
5.1.3. Main Results.
We summarize our results on the Amazon Reviews benchmarks in Table 1. Overall, models with sets of discrete item features outperform other competing methods that make use of single-embedding-based item features, while AgenticTagger helps achieve the best overall performance compared to various baselines, demonstrating its effectiveness.
Interestingly, the performance improvements are more significant in the Sports and CDs domains, with 8.9%-11.6% improvements over the closest competing model in NDCG@10, whereas the improvement is moderate (2%) in the Beauty domain. We hypothesize that this could be because AgenticTagger yields the most benefit when it is able to construct its vocabulary from a larger item corpus, where the natural language feedback aggregated via the Distill operation provides more stable signals. We further discuss this hypothesis in Section 6.2, where we show that AgenticTagger benefits from larger datasets with diverse items, even when doing so at the cost of pruning the vocabulary-building process.
Figure 3. Comparison against feature-crossing-based methods in NDCG@10. For ActionPiece, we compare against its ablated variant without inference time model ensembling, to isolate the effect on item-content understanding.
5.1.4. Comparisons with ID Crossing.
In light of recent advances in demonstrating the benefit of incorporating dataset collaborative filtering statistics directly into the tokenization process via feature crossing (Hou et al., 2025b; Singh et al., 2024), we also compare our method with two recent tokenization methods that merge frequently co-occurring semantic IDs into new IDs based on observed user behavior sequences, SentencePiece (Singh et al., 2024) and ActionPiece (Hou et al., 2025b). To be specific, these methods represent an alternative path towards better features for recommender systems: automated feature crossing based on user behavior. As shown in Figure 3, AgenticTagger remains a strong tokenization method, demonstrating the effectiveness of leveraging LLMs’ content understanding abilities in tokenization in addition to collaborative filtering signals. Meanwhile, we note that AgenticTagger is orthogonal to these crossing-based techniques, and we leave the crossing of AgenticTagger features to future study.
5.2. Ranking
Table 2. Ranking performance comparison between the Item ID baseline and AgenticTagger.
Metric Item ID (Baseline) AgenticTagger (Ours)
R@5 0.2655 $\pm$ 0.0037 0.2823 $\pm$ 0.0038
N@5 0.1966 $\pm$ 0.0029 0.1870 $\pm$ 0.0027
R@10 0.3456 $\pm$ 0.0040 0.4300 $\pm$ 0.0041
N@10 0.2224 $\pm$ 0.0029 0.2346 $\pm$ 0.0027
R@20 0.4427 $\pm$ 0.0042 0.6082 $\pm$ 0.0041
N@20 0.2470 $\pm$ 0.0029 0.2796 $\pm$ 0.0025
R@50 0.5761 $\pm$ 0.0041 0.8514 $\pm$ 0.0030
N@50 0.2733 $\pm$ 0.0027 0.3280 $\pm$ 0.0022
5.2.1. Dataset, Baseline, and Evaluation.
Besides generative recommendation, another frequent use case for structured item features is ranking (Singh et al., 2024; Zhu et al., 2021). To this end, we evaluate the effect of AgenticTagger features on a private dataset collected from a news-feed serving platform over a random subset of 47,358 users across an 8-day period. This results in 67,097 items and 206,345 interactions, where raw item content and user interactions are fully accessible. Following prior work in the literature (He et al., 2017; Kang and McAuley, 2018), we randomly sample 100 negative items per user to reduce computational overhead, and report performance with Recall@$K$ and NDCG@$K$, where $K\in\{5,10,20,50\}$, similar to our experiments on public benchmarks. The backbone recommendation model is a neural feature-aware model aligned with the deployed model in the platform, where we use either one-hot item IDs (i.e., one embedding per item) or AgenticTagger descriptors as IDs during training. For good scalability, we adopt a 3-layer setting for AgenticTagger, and build the vocabulary using 30,000 randomly subsampled items, resulting in 3 descriptors per item. We break item collisions using an extra digit of a collision-resolver token following standard practice in the semantic ID literature (Rajput et al., 2023; Hou et al., 2025a; He et al., 2025; Singh et al., 2024).
5.2.2. Overall Performance.
We report our results in Table 2; as shown, AgenticTagger brings significant improvements to ranking accuracy compared to raw item IDs, showcasing the effectiveness of AgenticTagger features. On the other hand, an interesting observation is that the item ID baseline shows slightly stronger performance for NDCG@5. We hypothesize that this is because when the dataset is sufficiently large for good convergence, as in our case, memorization of item-specific collaborative filtering signals via unique per-item embeddings helps fine-grained ranking among top candidates. This is aligned with prior findings in the literature (Kang et al., 2021), where compressing one-hot item representations into hashed embedding tables sometimes requires (albeit marginally) a performance trade-off. We discuss further verification of this idea in the appendix, where we show that in smaller datasets where interactions are sparser, AgenticTagger-generated features exhibit larger improvements compared to one-hot item representations (Section B.2).
5.2.3. Cardinality.
As discussed in the previous paragraph, AgenticTagger achieves strong performance with a much lower-cardinality vocabulary, utilizing only 2.7k unique descriptors compared to the full one-hot item table of over 67k item embeddings, achieving effective vocabulary compression. This demonstrates that AgenticTagger effectively captures the distinctive characteristics of each item within its automatically inferred descriptor vocabulary, validating the effectiveness of our proposed framework in enabling LLMs to understand and capture frequent patterns in the item corpus.
5.3. Interpretability and Controllability
Table 3. Performance improvements after single-turn critique on 5000 samples.
Dataset Metric Vanilla +Critique Improve.
Sports R@5 0.0318 0.0428 +34.59%
N@5 0.0206 0.0293 +42.09%
R@10 0.0500 0.0670 +34.00%
N@10 0.0264 0.0371 +40.31%
Beauty R@5 0.0504 0.0864 +71.43%
N@5 0.0350 0.0602 +72.17%
R@10 0.0738 0.1284 +73.98%
N@10 0.0425 0.0738 +73.65%
CDs R@5 0.0546 0.0666 +21.98%
N@5 0.0374 0.0445 +19.21%
R@10 0.0816 0.0990 +21.32%
N@10 0.0460 0.0549 +19.42%
As discussed earlier, a core advantage of LLM-generated features is their explainability, which naturally suits user-facing use cases, such as situations where one would like to control or constrain the matching mechanism of a recommender system. In this section, we discuss the value AgenticTagger features add to the interpretability and controllability of a recommender system via a single-turn critique setting, where a user gives feedback to the recommender system, providing information on the item they would like to interact with next. In particular, this setting is a fundamental feature in many conversational recommender systems (Li et al., 2018; Wang et al., 2022; He et al., 2023; Zhu et al., 2025b), and commonly manifests in production recommender systems as feedback with keyphrase-based rationales.
Inspired by recent work on using LLMs for user simulation in recommender systems (Bougie and Watanabe, 2025; Park, 2025; Zhang et al., 2025a; Wang et al., 2024f) and conversational recommendation (Wang et al., 2023; Zhu et al., 2025a), we make use of an LLM as a user simulator to evaluate whether AgenticTagger can enable effective user control over recommender system behavior. Specifically, we adopt a communication game in an information bottleneck setting, where an LLM (gemini-2.5-flash in practice) observes the descriptions of target items in its prompt, and is asked to select all potentially relevant top-level (i.e., level 1) features in the vocabulary, which generally contains a small number of keywords (e.g., 16 for the CDs and Vinyl domain, the largest public benchmark we evaluate on). We then generate the next item from a pre-trained generative recommender backbone using beam search, following standard practice, except we constrain its search space to be within the allowed level 1 features provided by the user simulator. While this interaction paradigm essentially “leaks” ground-truth information to the recommender, it verifies that (1) AgenticTagger produces a vocabulary that precisely identifies a meaningful subset of items, and (2) the produced vocabulary aligns well with the semantic understanding of LLMs, which are known to align well with human perception (Chang and Bergen, 2024), demonstrating the interpretability and controllability potential of AgenticTagger features.
As shown in Table 3, this constrained recommendation setting increases model performance significantly across datasets, showcasing that AgenticTagger features effectively communicate the nuances of items in the dataset and, importantly, in an interpretable manner, such that the behavior of the corresponding recommender system using AgenticTagger features can be straightforwardly controlled by natural-language-based critiques.
6. Ablations and Discussions
6.1. Effect of Multi-agent Self-refinement
Figure 4. Distribution of coverage changes between optimization steps across layers. We report level 2 to level 6, which have enough trials to obtain meaningful results for analysis.
In this ablation, we examine the effect of the iterative refinement loop in AgenticTagger on improving item coverage. To verify its effect, we plot the average percentage differences in item coverage (i.e., items that are assigned at least one feature among the current vocabulary) for each clustering job in Figure 4. As shown, the overall trend is that iterative refinement helps improve item coverage, leading to a reduced number of items that cannot be described by the vocabulary as the number of optimization steps grows. While we report the difference between steps 1 and 2 on the Sports dataset, the trend remains consistent across datasets and optimization steps.
On the other hand, we observe that iterative optimization tends to bring less steady coverage improvements to later-layer (i.e., finer-grained) AgenticTagger features. To this end, we hypothesize that to effectively improve item coverage, the LLM must be able to understand increasingly nuanced item-to-item similarities as the hierarchy depth increases; intuitively, it is easier to separate high-level differences, such as Rock music and Blues music, than specific subgenres of Rock music.
6.2. Scaling AgenticTagger to Larger Datasets
In AgenticTagger, a key decision to make when scaling to larger datasets is the balance between controlling the maximum branching factor $b$, i.e., how many features an item can be associated with at each granularity level, and how much data to subsample for generating the vocabulary using Algorithm 1. While these two techniques are not mutually exclusive, it is nevertheless helpful to understand which tends to yield better performance for a large dataset.
To investigate this trade-off, we compare the quality of AgenticTagger features generated from a random subset of 20k items with that of the original scaling choice we made on the CDs and Vinyl domain, where we constrained the branching factor to 1. That is, we only consider one randomly selected associated feature at each feature level for an item when generating finer-grained features in the next level. We found that, surprisingly, constraining the branching factor yields significantly superior performance compared to subsampling, where the performance drops to a level comparable to P5-CID. We hypothesize that constructing the vocabulary from more items helps AgenticTagger assign valid features to long-tail item groups, which are crucial for capturing niche user interests.
6.3. Comparison with Free-form Generation in Generative Recommendation
As discussed earlier, it is difficult to directly use free-form generated features for generative recommendation, as this method assigns sets of features to items with no constraints on vocabulary choice. To illustrate this point, we discuss an earlier attempt aimed at directly using free-form generated features as semantic IDs in the Sports domain.
We first prompt an LLM to generate free-form descriptors and treat all ever-generated descriptors as part of the valid vocabulary. We then train a dual-encoder model that predicts which descriptors are generated for each item among the possible descriptor vocabulary (i.e., every descriptor that has been generated by the LLM), following prior work in LLM-based document topic mining (Xie et al., 2025), where positive signals come from LLM generations. This process ensures that the potential match between each descriptor-item pair is sufficiently explored, even when the LLM did not generate a particular descriptor initially. Finally, we perform retrieval-augmented descriptor generation with another pass of prompting, except that the LLM this time is asked to select all relevant descriptors from the top 30 suggested descriptors predicted by the aforementioned dual-encoder model. This process yields a maximum of 27, a minimum of 0, and an average of 12.25 descriptors per item.
Due to the large number of descriptors per item, we then prune the descriptors assigned to each item via two methods. (1) Frequency: we split the tags by their occurrence across the corpus into four equal bins, ignoring tags with frequencies smaller than 10 or larger than 2,000, since these are overly rare or common. Each item’s descriptors are then selected by taking the most frequent tag from each popularity bin, where available, creating a coarse-to-fine tag sequence similar to semantic IDs. (2) K-means: we select 5,000 centroids by running K-means clustering over the text embeddings of the descriptor vocabulary, and swap the descriptors with their assigned centroid IDs, reducing the number of descriptors per item.
However, this framework results in poor overall performance: $\text{NDCG}@10$ for frequency-based pruning is 0.185, and for K-means-based pruning, it is 0.184, neither of which is comparable to modern semantic ID performance. This demonstrates, in contrast, the effectiveness of AgenticTagger in generating hierarchically structured, semantic-ID-like descriptors.
Table 4. Comparison on AgenticTagger performance with different LLMs on generative recommendation.
Backbone Sports Beauty CDs
N@5 N@10 N@5 N@10 N@5 N@10
Gemini-1.5 0.0184 0.0237 0.0319 0.0397 0.0349 0.0439
Gemini-2.5 0.0194 0.0251 0.0332 0.0408 0.0358 0.0453
6.4. Performance with Different LLMs
In our generative recommendation experiments on public benchmarks, we opted to use the gemini-2.5 family of models, as they were released closest to the time the experiments were conducted. However, to investigate if our framework is robust to the choice of base models, we also experimented with gemini-1.5-flash and gemini-1.5-pro as annotator and architect LLMs, respectively. In our experiments, we find both models yield competitive performance, while gemini-2.5 maintains a slight advantage across datasets, likely due to its stronger overall abilities in text generation and understanding.
6.5. Case Study
Table 5. Case Study: AgenticTagger and free-form generation as features. Item content rewritten to preserve privacy.
Item Context Free-form AgenticTagger
Ground Truth Skipper Limited, L0: Business & Finance
Transformer stock, Order Book Growth, L1: Stock Market
820% 3-yr returns Power Transmission. L2: Individual Performance
User History L0: Business & Finance
(1) Defence bidder, 7k% 5-yr returns (1) Defence PSU, DRDO Bidding. L1: Stock Market
(2) Industrial PSU, 200% 5-mo returns (2) PSU Stocks, Market Returns. L2: Individual Performance, Order Book Analysis
Finally, to understand the improvement brought by AgenticTagger compared to free-form generation, we took an example from our private dataset to analyze the effect of features from both frameworks on connecting the user’s recent interests to the ground-truth item. As shown, the free-form-generated features, while of generally high coherence, failed to generalize across items due to random surface phrase choices; in contrast, AgenticTagger successfully identified the user’s broad interest in the stock market and detailed interest in stock performance. While in both cases LLMs are effective at extracting general user interest in stocks, AgenticTagger is the only method that can organically connect user history to the ground-truth interacted item, demonstrating the value of vocabulary-constrained, structured descriptor generation.
7. Conclusion
In this work, we introduce AgenticTagger, a general framework for prompting LLMs to generate natural language item-descriptor representations for recommender systems. We showcase the wide applicability of AgenticTagger across various applications, including generative retrieval, ranking, and controllability-oriented, critique-based recommendation. We also discuss the properties of AgenticTagger, such as its complexity and feature cardinality. In the future, we plan to investigate methods for improving effectiveness by introducing AgenticTagger implementations that consider collaborative filtering signals during the feature generation phase, following the recent success of automated feature crossing.
Appendix A Appendix
A.1. Implementation Details
Table 6. Statistics of public datasets used. “Avg.” denotes the average sequence length.
Datasets #Users #Items #Interactions Avg.
Sports 35,598 18,357 260,739 8.32
Beauty 22,363 12,101 176,139 8.87
CDs 75,258 64,443 1,022,334 14.58
Table 7. Hyperparameter settings for each dataset.
Hyperparameter Sports Beauty CDs
learning_rate 0.005 0.005 0.01
n_semid 6+1 3+1 6+1
vocab_size 2487+80 1103+1145 742+192
warmup_steps 10,000 10,000 10,000
dropout_rate 0.1 0.1 0.1
weight_decay 0.07 0.10 0.07
beam_size 50 50 50
num_layers 4 4 4
d_model 128 128 256
d_ff 1,024 1,024 2,048
num_heads 6 6 6
d_kv 64 64 64
optimizer adamw adamw adamw
lr_scheduler cosine cosine cosine
train_batch_size 256 256 256
max_epochs 200 200 200
early_stop_patience 20 20 20
In this section, we cover our implementation details on public benchmark experiments disclosed in this work.
A.1.1. Baselines and pre-processing.
We directly reuse the data processing implementation open-sourced by Hou et al. 2025b for fair and straightforward comparison. We reuse the reported baseline performance in Hou et al. 2025b, where baseline results for BERTRec, SASRec, FDSA, and S³-Rec on the Sports and Beauty domains are taken directly from each respective work, and the rest are re-implementations by Hou et al..
A.1.2. Hyperparameter for GR Training.
We train the generative recommender model from scratch with AgenticTagger as semantic IDs for a maximum of 200 epochs, breaking the training if there are no improvements in $N@10$ for 20 epochs. Following Hou et al. 2025b, we set training batch size to 256 and tune learning rate from $\{1\times 10^{-3},3\times 10^{-3},5\times 10^{-3}\}$ and weight-decay from $\{0.07,0.1,0.15,0.2\}$. We conduct three repeated experiments with random seeds $\{2025,2026,2027\}$, and report the test set performance with the model checkpoints with best average $N@10$ across runs.
A.1.3. GR Inference.
We use standard auto-regressive decoding with a beam size of 20 for inference. However, since AgenticTagger can produce variable length semantic IDs, we do not sample from the base generative recommender for a fixed number of steps; instead, we generate until we reach the end-of-sequence token. The generated sequences are then ranked by their beam scores.
A.1.4. AgenticTagger
By default, we set the initial cluster size to 15 across all datasets and the number of refinement cycles to 3 (while the architect LLM can choose to exceed this number). The minimum cluster size for sub-clustering is set to 30, i.e., any group of more than 30 items would trigger the procedure for producing sub-category descriptors for the said group. We break the iterative refinement loop if the item coverage for any cluster is greater than 95 or if the refiner LLM fails to generate more than 20 error cases. Due to the large number of items in the CDs domain, we set the sub-clustering threshold to 100 for the branching-controlled version of the algorithm and 500 for the subsampled variant. While the final vocabulary size of AgenticTagger varies due to the stochasticity of program execution, we report the base vocabulary and the number of conflict resolvers (i.e., resolving tokens used when two items hash to the same set of AgenticTagger tags) as “base tokens + resolvers” in Table 7 under “vocab_size.” As shown, AgenticTagger is significantly more efficient than its semantic ID counterparts, using fewer than 3,000 words across the board on academic benchmarks.
Appendix B Further Related Works
ML Models with Natural Language Parameters. Beyond recommender systems, there have been continuing efforts to introduce parameters tied to explicit natural language concepts into ML models, e.g., concept-bottleneck models (Koh et al., 2020; Sun et al., 2025), typically for their improved interpretability (Zhong et al., 2024). However, this stream of work generally does not aim for improved accuracy and frequently assumes there exist sets of concepts (in our case, features) that are readily available (Koh et al., 2020). The most relevant work to ours is perhaps the label-free concept bottleneck model (Sun et al., 2025), which prompts an LLM with a task description for concepts. However, our aim is to infer a set of features from large sets of items for recommendation, the nature of which is typically difficult to infer from the task description itself, unlike domains such as image classification with a fixed set of labels.
LLM for Text Mining At Scale Another relevant stream of work to AgenticTagger, as discussed throughout the paper, consists of recent advances in leveraging LLMs for scalable text mining, typically manifesting as topic/latent theme mining (Xie et al., 2025; Pham et al., 2024) or taxonomy mining (Wan et al., 2024). However, these frameworks generally do not consider downstream applications and usually impose sequential dependencies on the interaction between LLMs and text corpora (Wan et al., 2024; Pham et al., 2024), whereas AgenticTagger uses parallelized operations, making scaling to recommendation-scale item corpora feasible.
LLM-generated Semantic IDs Concurrent to this work, we note there are also recent explorations of LLM-generated natural language semantic IDs, dubbed Term IDs (Zhang et al., 2026), where one prompts an LLM for free-form generated semantic IDs over a batch of similar items for adapting pre-trained LLMs for recommendation. To this end, Term IDs impose fewer constraints on the vocabulary itself, since a pre-trained LLM is able to generalize information across semantically similar Term IDs. On the other hand, AgenticTagger aims to develop a descriptor vocabulary with minimal constraints on the downstream model. To illustrate this point, AgenticTagger enables a 2-layer transformer model (the same as Tiger (Rajput et al., 2023)) to perform better than the 0.4B variant of LLM-based GR with Term IDs on the Sports and Beauty domains, showcasing the effectiveness of AgenticTagger.
B.1. Sample Prompts (Sports Domain)
Architect - Generate Inital Vocabulary You are an expert taxonomy architect. Your task is to create a set of fine-grained sub-categories.
The items you are categorizing all belong to the parent category: "{parent_rule_description}".
Your new categories MUST be more specific subdivisions of this parent category. Do not simply repeat the parent category’s name. For example, if the parent is "Firearm Accessories", your sub-categories should be "Optics", "Holsters", "Cleaning Kits", etc., not "Firearm Accessories" again.
Primary Goal: Create a set of ˜{n_target_rules} mutually exclusive sub-categories that logically partition the parent.
{context_prompt}
Input Data:
Below are diverse product examples. Use them to understand the product landscape you need to partition.
{sample_text}
Task & Strict Output Format:
Devise a set of categories. For each, provide a "name" and a "description" with "INCLUDES" and "EXCLUDES" clauses. Respond ONLY with a single JSON object: {"categories": [{...}]}.
Your JSON Response:
Annotator {context_prompt}
Above, I have provided the parent category and some product examples.
Your Goal:
Propose a list of exactly
{n_target_rules} sub-categories.
Constraint - Overlap:
Sub-categories must be mutually exclusive.
Constraint - Coverage:
Together, they should cover most common items found in "
{parent_rule_description}".
Constraint - Specificity:
Do NOT create a "Miscellaneous" or "Other" category. Every category must have a specific, meaningful name.
Constraint - Style:
Use professional, industry-standard terminology.
Output Format:
Your response must be a JSON object with the following structure:
"categories": [
"name": "Category Name",
"description": "Definition of what is included",
"includes": ["Example 1", "Example 2"],
"excludes": ["Non-example 1"]
Annotator - Generating Error Feedbacks You are a data analyst specializing in taxonomy quality control. Analyze a cluster of {len(ticket_cluster)} uncategorized products and synthesize a single, structured change proposal.
Existing Category Rules:
{existing_rules_text}
Analysis of Uncategorized Items:
{ticket_examples_text}
Task: Decide the most logical action and formulate a proposal. Respond ONLY with a single JSON object. Your response MUST strictly follow one of the formats below.
1. To create a new category:
{"change_type": "CREATE_NEW_CATEGORY", "problem_summary": "<summary>", "suggested_change": {"new_rule_description": "New Category Name: INCLUDES: ... EXCLUDES: ..."}}
2. To expand an existing category (THIS IS A VERY STRICT FORMAT):
The "refined_description" MUST be a single complete string, starting with a name, then "INCLUDES", then "EXCLUDES".
GOOD EXAMPLE:
"change_type": "EXPAND_EXISTING_CATEGORY",
"problem_summary": "The current ’Outdoor Gear’ rule is too generic and should explicitly include tactical accessories.",
"suggested_change": {
"rule_id_to_refine": "rule_a4368cef",
"refined_description": "Outdoor & Tactical Gear: INCLUDES: Tents, backpacks, sleeping bags, and tactical accessories like gloves, belts, and pouches. EXCLUDES: Specialized sporting equipment, firearms, and knives."
3. To ignore outliers:
{"change_type": "IGNORE_AS_OUTLIERS", "problem_summary": "<summary>", "suggested_change": {"reason": "<reason>"}}
Your JSON Response:
Architect: Taxonomy Refinement You are a senior taxonomy manager. Review the following change proposals and decide whether to approve or reject each. Your goal is a clean, non-overlapping taxonomy. But if the proposal is reasonable, you should try to accomodate that. You should also reject proposals that are out-of-place by common sense. Respond ONLY with a JSON list of objects, each with "proposal_id", "decision" (’APPROVED’ or ’REJECTED’), and "reasoning".
Proposals for Review:
{proposals_text}
Your JSON Decision List:
B.2. Ranking Performance on Smaller Dataset
Table 8. Ranking performance comparison between Item ID baseline and AgenticTagger.
Metric Item ID (Baseline) AgenticTagger (Ours)
R@5 0.0833 $\pm$ 0.0068 0.1205 $\pm$ 0.0080
N@5 0.0589 $\pm$ 0.0051 0.0853 $\pm$ 0.0060
R@10 0.1163 $\pm$ 0.0079 0.1751 $\pm$ 0.0093
N@10 0.0695 $\pm$ 0.0052 0.1028 $\pm$ 0.0061
In this section, we discuss further performance comparisons in ranking experiments, where we report both Recall and NDCG in Table 8. This subset contains 1,791 users, 15,830 items, and 47,260 interactions, making the collaborative filtering signal less salient due to the smaller dataset size. As shown, our method consistently outperforms the ID-based method by a larger margin than in the main experiment, verifying our intuition: when the dataset is large and the collaborative filtering signal is more stable, assigning more item-specific parameters helps in memorizing item-specific information, sometimes leading to better ranking performance. Conversely, when the dataset is sparse, content-aware structured discrete features shared among items help convergence significantly, making AgenticTagger features potentially applicable to cold-start settings, where they lead to better generalization.
B.3. Cardinality of Free-form Generation
To investigate whether free-form generation indeed leads to vocabulary explosion, we conduct an ablation study where we prompt an LLM to generate three keywords over the private dataset discussed in Section 5. Indeed, there were approximately 90k unique keywords generated over 210k documents; however, only 21k (or 24%) of the descriptors appeared in more than one item. Consequently, these features are inapplicable for natively acting as compact, ID-like features, making downstream modeling challenging. On the other hand, descriptors generated by AgenticTagger have a much higher utilization rate (87%) across a vocabulary of over 2k, demonstrating their effectiveness.
B.4. Tags Generated
We show qualitative examples of the discovered vocabulary in Table 9. For presentation purposes, we pick two descriptors per level but only expand on one of them until no valid children are found. Note that in AgenticTagger, the depth of each hierarchy path can be of variable length, as shown in the CDs domain example. As showcased in the table, AgenticTagger is able to find clear separations of products, especially towards the earlier layers of the hierarchy (e.g., “Skincare” and “Haircare” products). However, towards the end of the hierarchy, the boundary between products becomes fuzzy, and it is more challenging to come up with clean, disentangled separation rules (e.g., “Specialty Function” and “Ceramic Plate Irons”), making it harder to enforce a nominal constraint on features. While we partially address this issue by instructing the LLM to select the best-fitting descriptor for each product when multiple are technically correct, developing methods to fundamentally address such disentanglement issues remains an open problem. Nevertheless, even at later layers, AgenticTagger can still discover meaningful descriptors that separate the products, such as sleeping bags by their applicable temperature range, as shown in the Sports domain.
B.5. Effect of Multi-agent Self-refinement on Other Domains
We showcase the improvements in coverage by reporting the distribution of delta changes in the coverage rate in Figure 5. The trend is consistent across domains: (1) our designed multi-agent self-refinement loop effectively refines the vocabulary, improving item coverage; and (2) the trend of improvement (i.e., delta greater than zero) becomes less stable in later layers, possibly because it is more challenging to generate disentangled descriptors, as qualitatively analyzed in Section B.4. Regardless, the Wilcoxon signed-rank test shows that improvements are significant (i.e., delta greater than zero) with p-values smaller than 0.05, except for levels five and six in the CDs domain, corroborating the general effectiveness of the self-refinement mechanism.
Figure 5. Coverage improvements across domains. We discuss further details in Section B.5.
Table 9. Discovered Vocabulary (Sub-sampled) Across Domains
Beauty Domain Music (CDs) Domain Sports Domain
L1: Skincare - Facial Treatments L1: Jazz Music L1: Fitness & Exercise Equipment
L1: Haircare - Treatments & Styling L1: Electronic Dance Music (EDM) L1: Camping & Hiking Gear
   L2: Hair & Scalp Treatments    L2: Smooth Jazz    L2: Shelter Systems
   L2: Hair Styling Tools & Acc.    L2: Free Jazz & Avant-Garde    L2: Sleeping Systems
      L3: Hair Dryers       L3: Interdisciplinary Avant-Garde       L3: Sleeping Bag Liners
      L3: Flat Irons & Straighteners       L3: Free Improv. & Avant-Garde       L3: Sleeping Bags
         L4: Travel & Mini Flat Irons          L4: Free Jazz (Post-Bebop)          L4: Sleeping Bag Shapes
         L4: Professional Salon Irons          L4: Avant-Garde (Exp.)          L4: Temp.-Rated Bags
            L5: Specialty Function             L5: Atonal Exploration             L5: Mild to Cool Bags
            L5: Ceramic Plate Irons             L5: Genre-Infused Exp.             L5: Cold Weather Bags
               L6: Adv. Plate Tech.                L6: 10-30 Degree Bags
               L6: Ionic Technology                L6: 0-Degree Rated Bags
Table 10. Summary of Notations
Notation Description
$\mathcal{U},\mathcal{I}$ Sets of users and items
$S_{u}$ Interaction history for user $u\in\mathcal{U}$
$X_{i}$ Raw attributes for item $i\in\mathcal{I}$
$\Pi,\pi$ LLMs, and a specific instance
$\mathcal{E}$ Set of natural language error feedback
AA Descriptor feature vocabulary
$a$ A specific feature descriptor ($a\in A$)
$g(\cdot)$ Transformation mapping attributes to a subset of AA
$f_{\theta}$ Recommender system parameterized by $\theta$
Table 11. Impact of the backbone model used in AgenticTagger on downstream performance, Recall and NDCG.
Backbone LLM Sports Beauty CDs
R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10
Gemini-1.5 0.0284 0.0184 0.0448 0.0237 0.0469 0.0319 0.0710 0.0397 0.0531 0.0349 0.0812 0.0439
Gemini-2.5 0.0299 0.0194 0.0474 0.0251 0.0492 0.0332 0.0728 0.0408 0.0544 0.0358 0.0836 0.0453